A Complete Bibliography of Publications in Communications in Mathematics and Statistics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 November 2024 Version 1.31

Title word cross-reference

(α, β) [115]. ∗ [85]. 1 [45]. {18, 15, 9; 1, 1, 10} [78]. 1: 1 [36]. 2 [50, 83, 138, 204]. 2^{4} [138]. 3 [132, 77, 23, 18, 196, 134, 4, 15]. α [36]. β [36]. C^{0} [8]. C^{1} [216]. c_{2} = 2 [257]. CP_{2} [127]. D [252]. f [292, 223]. f(x) = x^{3} + x [323]. F_{2} + uF_{2} + u^{2}F_{2} [202]. G [161, 305, 90]. GARCH(p, q) [219]. H [139]. k [317, 277]. L [46, 20]. L^{\frac{3}{2}} [23]. L^{n} [221]. λ_{0}(g_{0}) = 0 [49]. log [213]. M [207, 103, 100]. R^{8} [51]. H [108]. H(div) [2]. X [142]. CP [99]. P [99]. μ [314]. N [30, 103, 155]. p [211, 27, 121, 255, 58]. p^{3} [58]. ∂ ¯∂ [297]. π_{1} [91]. R [42]. SαS [210]. σ [316, 247]. T [70, 230, 212]. U [20]. U_{r,s}(}↕_{n}) [42]. F [153]. X [131]. x^{−s} [84]. x^{−s} ln |x| [84]. x_{+}^{μ} [84]. |x|^{μ} [84].

-Absorbing [83]. -Adic [255]. -Ary [103]. -Bases [314]. -Condense [99]. -Condensing [99]. -Conforming [2]. -Embedded [100]. -Estimate [8]. -Estimation [207]. -Expectation [90]. -Generalized [131]. -Groups [211, 138, 58, 153, 230]. -Harmonic [292, 27]. -Hypercentral [223]. -Lemma [297]. -Lie [85]. -Manifold [23]. -Manifolds [77, 18, 161]. -Matrix [42]. -Nearest [277]. -Neutral [103]. -Norm [23]. -Optimal [252]. -Permutability [247]. -Permutable [108]. -process [212]. -Regular [196]. -SDEs [305]. -Solubility [316]. -Spheres [4]. -splines [70]. -Statistics [20]. -Step [132]. -Subgroups [204, 142]. -Supersolvability [121]. -th [317]. -Transforms [30].

3D [313].

A-type [318]. Abelian [58, 204]. Absorbing [83]. Academic [95]. Accelerated [132]. Accessing [95]. Accurate [312]. Across [224]. Acting [82]. Actions [109, 223]. Adaptive [176]. Added [173]. Additive [299, 102, 185]. Adhesion [294]. Adic [255]. Affine [318]. Air [135]. ALE [179]. Algebra [175, 237]. Algebraic [86]. Algebras [82, 85, 266]. Algorithm [128, 314, 134]. Almost [117, 137, 110]. Along [49, 54]. Alternating [55, 107]. Alternative [170, 320]. Always [184]. Among [227, 215]. Analysis [95, 248, 132, 88, 319, 309, 290, 182, 180, 225, 39, 263, 53, 70]. Analysis-Suitable [70]. Analytic [255]. Anderson [280]. Application [60, 115, 112, 148, 61]. Applications [190, 152, 71, 79, 242, 244, 283, 201, 159]. Approach [71, 79, 178, 290, 38, 158]. Approaches [9]. Approximation [319, 34]. Approximations [238]. Arbitrary [70]. Area [287, 135]. Area-Preserving [287]. Arithmetic [18, 68, 215]. Array [78]. Articles [28]. Artificial [309]. Ary [103]. Associated [302, 160, 283, 228]. Asymmetric [219, 213]. Attribute [115]. Automorphism [76]. Automorphisms [78]. Auxiliary [271, 171, 75, 126]. Average [193]. Axiom [253]. Axis [254].

B [134, 238]. B-Spline [134, 238]. Backward [74, 122, 111, 60, 191]. Ball [254]. Banach [82]. Barrier [303]. Base [256]. Based [195, 188, 122, 128, 264, 93, 36, 116, 269, 302, 273, 101, 225, 231, 260, 173, 263, 246, 254]. Bases [2, 314, 123]. Bathtub [102]. Bathtub-Shaped [102]. Bayesian [242, 172, 165, 159, 212, 158]. Beltrami [19]. Benford [184]. Bergman [181, 3]. Berry [129, 20]. Besov [245]. Between [17, 114]. BGK [261]. BI [262]. BI-GreenNet [262]. Big [173]. Biharmonic [289]. Bilinear [218]. Binary [155]. Binomial [234]. Biquadratic [113]. Birkhoff [200]. Bladder [302]. Blocked [136]. Blossoming [193]. Blow [130]. Blow-Up [130]. Bogomolov [236]. Bogomolov-Type [236]. Boltzmann [105, 261]. Bonnet [282]. Bott [297]. Boundary [118, 88, 48, 262, 15, 263, 130]. Boundary-Mapping [88]. Bounded [292, 198]. Bounded-Gap-Hitting [198]. Bounds [129, 20, 301]. Boussinesq [50]. Braces [239]. Breiman [152]. Brownian [267, 282, 6, 228, 150]. Bundles [236, 255, 94]. BV [237].

Cahn [312]. Calabi [72, 5, 298]. Calibration [189]. Camassa [144]. Cancer [302]. Canonical [186]. Case [49]. Cauchy [321]. Causal [320]. Cell [312]. Censored [195, 188, 242, 172]. Censoring [132, 290, 112, 167]. Central [308]. Centralizers [131]. Centre [279]. Certain [298]. Chain [271]. Chains [243]. Change [114]. Chaos [233]. Chaotic [67]. Characteristic [283]. Characteristics [51, 135]. Characterization [163, 245]. Chen [201]. Chern [282, 297]. Chung [61]. Cities [224]. Class [105, 189, 318, 75, 149, 16]. Classification [211, 191]. Classifier [273, 277]. Classifying [159]. Clipping [93]. Closed [89, 194, 51, 67]. Closed-Form [194]. Closure [261]. Clouds [34]. Cluster [176, 225]. CNC [254]. Cochain [237]. Cocycles [253]. Codes [202]. Coding [36]. Coefficient [97]. coefficients [297]. Cohomology [18, 297]. Colored [147]. Columns [182]. Combinatorial [322, 13, 163, 222]. Combinatorics [199].

Commodity [33]. Common [215]. Commutative [311, 83]. Commutator [77]. Commutators [226]. Compact [51, 3, 186, 91, 162, 106]. Complementary [195]. Complete [242, 99, 26, 32]. Complex [243, 18, 237]. Complexity [157]. Component [172]. Composed [24]. Composite [270]. Compositions [84]. Compound [136]. Compression [134]. Compute [314]. Concavity [27]. Concerning [28, 125]. Condense [99]. Condensed [37]. Condensing [99]. Condition [284, 200, 130]. Conditions [288, 243, 25, 15]. Confined [312, 296]. Conformal [300]. Conforming [2]. Congested [135]. Congruences [196]. Conics [32]. Conjecture [140, 206]. Connected [224, 296]. Consequences [256]. Conserved [294]. Consistency [277]. Constraint [178]. Constraints [306, 107]. Construction [72, 119]. Contacts [7, 14, 21, 29, 35, 41, 59]. Context [217]. Continuity [10]. Continuous [218]. Continuous-Time [218]. Continuum [166]. Control [74, 71, 79, 288, 25, 60, 169, 160]. Controlling [229]. Convergence [64, 149, 324]. Convex [24, 51, 27, 197]. Convolution [321]. Coordinates [179]. Copolymer [312]. Copula [227, 203]. Correlations [310]. Correspondence [255, 263]. Coupled [74]. Coupling [6, 5]. Covariance [165, 136]. Covariations [210]. Criteria [316, 45, 121, 153]. Criterion [300]. Critical [46, 1, 168]. Cubic [216, 169]. Current [290]. Curvature [23, 143, 27, 275]. Curve [313]. Curves [140, 169, 34]. Curves/ [34]. Customer [154]. Cutters [254]. Cycles [86]. Cyclic [237, 202]. Cylinder [24].

D [45, 50, 134, 15]. Data [195, 188, 276, 271, 50, 302, 290, 273, 165, 300, 112, 201, 158, 173, 277, 241]. Data-Based [173]. Dates [28]. Decision [115]. Decision-Making [115]. Decomposable [66]. Decomposition [278]. Decoupled [312]. Deep [122, 128, 259, 185, 231, 263, 246]. Defined [174]. Defining [84]. Deformation [264]. Degree [272, 70]. Delay [160]. Delta [71, 79]. Density [273, 234]. Dependence [272]. Dependent [152, 224, 277]. Derivable [85]. Derivations [73, 131, 82]. Derivatives [125]. Design [169]. Designing [250]. Designs [252]. Detecting [17]. Determine [98]. Determining [97]. Deviation [90]. Deviations [151]. Diagnostic [315]. Diagonal [3, 98, 148]. Diameter [257]. Diamond [192]. Diblock [312]. Dichotomous [170]. Dictionary [55]. Diffeomorphic [313]. Diffeomorphisms [307]. Difference [125, 30, 63, 245, 215]. Different [190, 68]. Differentiable [264]. Differential [122, 19, 226]. Diffusion [47, 303, 150]. Diffusions [288]. Dimension [266]. Dimensional [33, 248, 122, 229, 144, 151]. Dimensions [260]. Dirac [38]. Direction [55, 107]. Discontinuous [24, 144, 53]. Discovery [229]. Discrete [45, 294, 303, 243, 164]. Discrete-Time [243]. Discretizing [19]. Discriminator [323]. Disease [232]. Distance [292, 78, 257]. Distance-Regular [257]. Distinct [196]. Distinguish [114]. Distortion [235]. Distributed [306]. Distribution [190, 195, 188, 210, 242, 116, 302, 183, 101, 112, 182, 283, 201, 220]. Distributions [208, 214, 52]. Divergence [2, 11, 244, 220, 234]. Divergence-Free [2, 11]. Divisor [311]. Divisors [12]. DNA [202]. does [98]. Domain [218, 189]. Domains [319, 50, 180, 187, 174]. Donsker [79, 71, 61]. Double [56, 126, 167]. Doubling [72]. Doubly [74, 172, 320]. Drift [267, 228]. Drifts [47, 293]. Driven [111, 151]. Dropout [241]. Drops [112]. Duration [111]. Dynamical [104, 322, 222].

Each [138]. Eccentric [182]. Effect [95]. Effects [232]. Efficient [294, 189, 171, 75]. Eigenvalue [240, 180]. Eigenvalues [89, 243]. Eight [28]. Einstein [8, 145]. Elastic [24]. Elegant [184]. Element [11]. Elliptic [324, 228, 53]. Elliptical [214]. Embedded [100]. Embedding [243, 98]. Empirical [101]. End [254]. Energy [22]. Entire [63, 221]. Entropy [307, 115, 13, 269, 304, 110]. Equal [133, 161]. Equation [280, 289, 261, 144, 97, 197, 26, 130]. Equations [47, 16, 105, 122, 50, 30, 187, 168, 260, 15, 228, 53]. Equivariant [161]. Ergodic [109]. Erratum [28, 79]. Error [319, 229, 309, 300, 296]. Error-in-Variables [229]. Errors [252, 212]. Esseen [129, 20]. Estimate [143, 97, 8]. Estimates [47, 319, 89, 148, 26]. Estimating [271, 310, 167]. Estimation [177, 190, 195, 188, 218, 219, 267, 207, 299, 242, 126, 317, 112, 212, 320, 241, 249]. Estimations [301]. Estimator [117, 137, 170, 171, 234]. Estimators [189, 133, 75]. Euclidean [6, 221]. Euler [50]. European [194]. Evolving [209]. Example [302]. Exciting [291]. Exclusion [146]. Existence [141, 205, 324]. Expansion [233, 3]. Expansions [179]. Expansive [67]. Expectation [272, 61, 90]. Expectations [283]. Expectile [317]. Experimental [184]. Experiments [250]. Explicit [216, 279]. Exponential [195, 207, 115, 133, 126, 116, 112, 301, 167, 130]. Exponentiated [242, 116]. Extended [87, 318, 212]. Extension [201, 110]. Extensions [73, 87, 152, 217, 200, 206].

Face [109, 185]. Factor [276, 248, 194]. Factorised [239]. Factors [48]. False [229]. Families [321, 275, 196, 5, 91]. Family [208, 209]. Fano [72]. Features [65]. Fermat [296]. Fibonacci [199]. Fibonacci-Like [199]. Field [312, 294, 25, 60, 160]. Fields [184, 322, 149, 113]. Filtered [278]. Filtrations [186]. Finance [71, 79]. Finding [93]. Finite [316, 211, 265, 48, 66, 108, 142, 11, 251, 223, 121, 295, 92, 75, 138, 58, 100, 124, 230, 204, 110]. First [89]. Fit [101]. Fitting [238]. Five [254]. Five-Axis [254]. Flat [179]. Flexible [188, 276, 242]. Floer [322, 164]. Flow [118, 89, 23, 157, 143, 253, 64, 4, 49, 40, 54]. Flows [209]. Fluids [151]. Fokker [16]. Forcibly [177]. Forecast [173]. Form [194]. Forman [322]. Forms [37, 155]. Formula [307]. Formulae [194]. Formulas [256]. Forward [33, 74, 111, 60, 191]. Forward-Backward [111]. Founded [184]. Four [250]. Four-Level [250]. Fourfolds [72]. Fourier [226]. Fractional [280, 30, 301, 94, 272, 267]. Fractional-Degree [272]. Free [2, 166, 321, 11]. Frequency [218, 310]. Frequency-Domain [218]. Frobenius [318]. Fully [312, 74, 294, 141]. Fully-Decoupled [312]. Function [69, 52, 283, 102, 225]. Functional [71, 79, 46, 232, 308, 315]. Functionals [149]. Functions [262, 27, 125, 63, 227]. Fundamental [96]. Fusion [178]. Fuzzy [310, 115, 244, 225].

Galerkin [294, 24, 144, 53]. Gamma [268, 154]. Gamma/Weibull [154]. Gap [240, 198]. GARCH [213, 234]. Gauduchon [162]. Gauge [118]. Gauss [282]. Gaussian [216, 280, 27, 147, 149]. General [171]. Generalisation [230]. Generalization [247]. Generalized [190, 131, 82, 299, 322, 242, 20, 69, 164, 36, 302, 273, 64, 112, 220, 15]. Generators [279]. Genus [67, 113]. Geometric [195, 209, 263]. Global [298]. Gompertz [190]. Goodness [101]. Goodness-of-Fit [101]. Governed [160]. Grade [151]. graded [217]. Gradient [233, 26]. Graph [264, 240, 78, 156]. Graph-Based [264]. Graphs [286, 289, 251, 311, 257, 120, 324]. Green [262]. GreenNet [262]. Grids [11]. Group [279, 109, 42, 223, 121, 98, 285, 124]. Groups [239, 316, 77, 211, 103, 265, 48, 66, 108, 142, 251, 318, 76, 295, 127, 92, 153, 138, 58, 100, 230, 204]. Gruenberg [251]. Guarantee [174].

Half [181]. Handling [276]. Harmonic [179, 292, 27]. Harnack [26, 305]. Having [301]. Hazard [102]. Hazards [9]. Heat [197, 148]. Height [27, 296]. Hele [312]. Help [7, 14, 21, 29, 35, 41, 59]. Hermitian [143, 186]. Heteroscedastic [252]. Hierarchical [2, 252, 274]. Higgs [236]. High [248, 122, 229, 260]. High-Dimensional [248, 122, 229]. Higher [94]. Higher-Order [94]. Hilbert [113]. Hilliard [312]. Hitting [198]. Hodge [86]. Hodograph [169]. Hölder [47, 10]. Holm [144]. Holomorphic [186, 162]. Homeomorphisms [67]. HOMFLY [69]. Homogeneous [245]. Homology [322, 77]. Homotopy [77, 237]. Horizon [160]. Hörmander [96]. Horseshoe [198]. Hunt [139]. Hybrid [112]. Hydrodynamic [146]. Hyperbolic [307, 18, 97]. Hypercentral [223]. Hyperkähler [17]. Hypersurfaces [51]. Hypothesis [139, 136].

Ideals [83]. Identification [299]. Identities [103]. Identity [283]. II [318, 195, 188, 242, 112]. Imaginary [113]. Imbalanced [273]. Impact [24]. Impact-Induced [24]. Implicit [319, 180, 134]. Implicitly [174]. Improved [225]. Imputation [271]. Inaccuracy [36]. INAR [291]. Inclusion [192]. Incompressible [187]. Incorporating [178, 300]. Incorrect [28]. Independence [13, 136, 70]. Independent [139, 212]. Index [9, 58]. India [135]. Individual [95]. Induced [24]. Industrial [173].

Inequalities [301, 305, 40, 54]. Inequality [236, 26, 49, 221, 106]. Inference [116, 320, 291]. Inferences [302]. Infinite [33, 160, 196]. Infinite-Dimensional [33]. Infinity [179]. Information [244, 178, 220]. Informative [290]. Inhomogeneous [187]. Initial [238]. Initializations [309]. Insider [71, 79]. Integer [234]. Integer-Valued [234]. Integral [262, 301]. Integration [174]. Intensity [303]. Inter [114]. Inter-Onset [114]. Intermittent [241]. Internal [120]. Intersection [286, 78]. Intersections [32]. Interval [114, 227]. Invariance [261, 61]. Invariant [295]. Inverse [242, 99, 156, 97]. Inverted [112]. Inviscid [15]. Involving [226, 221]. Iso [263]. Iso-Geometric [263]. Isogeometric [88, 319, 180, 53, 246]. isometric [292, 161]. Isomorphic [251]. Isoparametric [275]. Isospectral [31, 265]. Itô [233]. Iterated [61, 90].

Jacobi [226]. Jensen [244]. Jet [94]. Join [99]. Join-Complete [99]. Joint [165]. Jointly [210]. Jump [303, 288]. Jumps [60].

Kähler [46, 64, 3, 8, 145]. Katok [307]. Kawaski [312]. Kegel [251]. Kernel [321, 3, 148]. Kernels [301]. Knot [216]. Köthe [206]. Krätzel [52].

Lagrangian [57, 168]. Landmarks [313]. Laplace [101, 19]. Laplacian [240]. Large [281, 250]. Largest [240, 268]. Latent [310]. Lattice [105]. Lattices [99]. Law [184, 61]. Laws [90]. Learning [276, 122, 104, 128, 259, 55, 261, 262, 231, 263, 246]. Learning-Based [122, 128, 231]. Lectures [86]. Left [239]. Lemma [297]. Lemmas [68]. Length [5]. Level [250, 27]. Lévy [111, 139, 160, 151]. Lie [85]. Life [132]. Lifetime [201, 154]. Like [199]. Likelihood [9, 101]. Likely [215]. Limit [146, 84, 15, 308]. Lindley [182]. Line [255].

Linear [117, 137, 310, 299, 171, 252, 70, 315, 61]. Link [77]. Liouville [22, 130]. Lipid [294]. Liu [117, 137]. Livsic [253]. Lizorkin [245]. Loading [182]. Local [9, 205, 24, 38, 144, 266, 297]. Location [224]. Location-Dependent [224]. Locus [256]. Loewner [141]. Log [190]. Logarithm [61, 90]. Logarithmic [233, 49, 54]. Logistic [190, 183]. Lomax [132]. Long [135]. Long-Term [135]. Longitudinal [165, 158, 241]. Look [172]. Low [248]. Low-Dimensional [248]. Lower [116].

Machine [104, 191]. Machining [254]. Macroeconomic [173]. Magnetohydrodynamics [2]. Making [115]. Manifold [23, 119]. Manifolds [17, 77, 18, 3, 318, 161, 186, 5, 162, 26, 249]. Mapping [88, 274]. Mappings [85]. Maps [17, 292, 199, 145]. Markets [33]. Markov [219, 243]. Marshall [208]. Martingale [308]. Martingales [129, 160]. Material [24]. Mathematical [44]. Matrices [37]. Matrix [243, 42, 253, 283]. Maximal [6, 295, 5, 110]. Maximum [60]. McCoy [200]. McKean [293]. Mean [271, 81, 189, 171, 25, 60, 126, 165, 160, 75]. Mean-Covariance [165]. Mean-Field [25, 60, 160]. Means [177]. Measure [235, 244, 36]. Measures [315]. Medium [260]. Melt [312]. Memory [306]. Meromorphic [125]. Mesh [65, 231, 119]. Meshes [281, 274]. Metahamiltonian [211]. Method [281, 105, 128, 62, 24, 93, 169, 55, 159, 107, 254]. Methods [190, 271, 122, 144, 112, 260, 53]. Metric [307]. Metrics [179, 186, 8, 145]. MHD [2]. Micro [112]. Micro-Drops [112]. Mills [143]. Minimal [66, 5]. Minimality [56]. Minimum [224, 234]. Missing [276, 271, 299, 291, 315]. Mixed [267, 34]. Mixing [157, 249]. Mixture [81, 172].

Model

[177, 117, 137, 276, 312, 132, 294, 9, 303, 172, 232, 261, 273, 102, 227, 225, 320, 154, 315]. Modeling [232, 165, 114, 203]. Models [33, 207, 299, 213, 194, 109, 229, 55, 65, 252, 212, 270, 234, 151]. Moderate [90, 151]. Modified [117, 137, 123]. Modules [87, 175, 284, 266]. Modulo [155]. Moment [116, 261, 145]. Monotonicity [89]. Morse [322, 164]. Moser [221, 106]. Most [138, 215]. Motion [267, 282, 228]. Motions [6]. Moving [34]. Multi [294, 126, 75]. Multi-Auxiliary [75, 126]. Multicategory [191]. Multidimensional [222]. Multilevel [95]. Multilinear [131, 82]. Multiobjective [94]. Multiple [34, 289, 115, 268]. Multiple-Outlier [268]. Multiplication [237]. Multiplicative [321, 301, 222, 305]. Multipliers [55]. Multivariate [248, 235, 290].

Nabla [30]. Navier [205, 187, 15]. Near [74]. Near-Relaxed [74]. Nearest [277]. Necessary [288, 25]. Negative [176, 234]. Neighbor [277]. Nested [285, 227]. Nets [156]. Network [262, 260]. Network-Based [260]. Networks [309]. Neumann [85, 168, 130]. Neural [264, 309, 260]. Neutral [103]. Neutrix [84]. Nirenberg [141]. Noetherian [284]. Noise [280, 16, 147, 305, 151]. Non [216, 16, 242, 264, 171, 66, 20, 199, 24, 251, 161, 266, 227]. Non-Convex [24]. Non-isometric [161]. Non-nested [227]. Non-projective [266]. Non-renormalizable [199]. Non-rigid [264]. Non-sensitive [171]. Non-Solvable [251]. Non-supersolvable [66]. Non-trace-class [16]. Non-uniform [216, 20]. Nonignorable [299, 241, 315]. Nonlinear [105, 141, 289, 160, 193, 26, 149, 324, 158]. Nonlocal [47, 96]. Nonmonotonic [232]. Nonparametric [249]. Nonresponse [241]. Nonsymmetric [237]. Norm [23, 221].

Normal [81, 66, 220, 283]. Normalized [129, 240, 237, 90]. Normalizer [204]. Note [56, 143, 236, 127, 43, 278]. Notes [207]. Number [286, 138]. Numerical [105, 122, 128, 180, 174].

Obata [43]. Observables [38]. Octonion [39]. Odd [190, 196]. Off [3]. Off-Diagonal [3]. Ohta [312]. Olkin [208]. On-diagonal [148]. One [314, 125, 110]. Onset [114]. Oort [140]. Open [217]. Operad [237]. Operator [125, 19]. Operators [45, 31, 99, 181, 226, 209, 96]. Optimal [71, 79, 250, 288, 25, 60, 252, 296, 160, 227]. Optimization [250, 246, 254]. Optimum [132]. Options [303, 194]. Order [312, 81, 288, 115, 36, 94, 138, 268]. Order- [115]. Ordering [208, 268]. Orderings [283]. Ordinal [158]. Ore [87]. Orientable [67]. Orientation [254]. Ornstein [267]. Outlier [268]. Overall [309].

Package [203]. Pair [37, 107]. Pair-Wise [107]. Pairs [162]. Pairwise [178]. Parabolic [280, 122, 62, 236, 26]. Parallelepiped [109]. Parameter [42, 167, 185]. Parameterization [287]. Parametric [314]. Parametrization [88]. Parametrizations [256]. Partial [299, 122, 8]. Partially [307, 92]. Partition [120, 70]. Partitions [196]. Parts [196]. Path [296]. Patodi [44]. Pattern [110]. Paw [251]. PBW} [217, 73, 206]. PDE [62]. Peng [308]. Percolation [166]. Performance [95]. Periodic [213]. Perks [102]. Permutability [247]. Permutable [108]. Perturbation [62]. Phase [312, 294, 75, 145, 168]. Phase-Field [312]. PHT [123]. PHT-Splines [123]. Pitch [114]. Plan [132]. Planar [50]. Planarity [163]. Planck [16]. Planning [296]. Poincaré [200, 1]. Point [256, 34, 97, 1, 313]. Pointwise [288]. Pollutants [135]. Polygon [169].

Polygonal [50]. Polyhedral [119].

Polynomial [69, 93, 323]. Polynomials [131, 82, 1, 63]. Population [271, 171, 75]. Populations [159]. Porous [260]. Portfolio [235, 60]. Positive [181, 300, 67]. Potential [294]. Power [302, 317, 234]. Prediction [188, 300]. Preface [258]. Presence [271]. Preserving [287, 261]. Price [33]. Pricing [303]. Primary [83, 100]. Prime [131, 82]. Primes [215]. Principle [60, 61]. Prior [178]. Probability [133]. Problem [74, 141, 60, 180, 97, 94, 225, 168]. Problems [128, 115, 217, 156]. Procedure [285]. Process [267, 111, 270, 212]. Processes [218, 219, 146, 233, 139, 160, 291, 150, 249]. Processing [231]. Product [152, 66, 226, 124]. Progressions [68, 215]. Progressive [195, 188, 132]. Progressively [112]. projective [266]. Pronormality [142]. Properties [190, 80, 133, 222]. Property [197, 204]. Proportion [229]. Proposal [104]. Pseudo [226]. Pseudo-differential [226]. Pythagorean [169]. Pythagorean-Hodograph [169]. Python [203]. pyvine [203].

QFT [38]. QML [219]. QMLE [213]. Quadratic [155, 260]. Quadrature [216]. Quantile [269, 158]. Quantile-Based [269]. Quantitative [171]. Quantities [19]. Quantization [130]. Quantum [279, 42]. Quasi [292, 153, 83]. Quasi- [153]. Quasi-isometric [292]. Quasiconcavity [197]. Quaternionic [17]. Quesney [237].

Radicals [206]. Ragas [114]. Ranchi [135]. Random [95, 152, 286, 309, 149]. Randomized [177, 170]. Rank [140, 91]. Rank-2 [91]. Rate [114, 102]. Ratio [189, 133, 273, 101]. Ratio-Based [101]. Rational [314]. Rayleigh [172]. Re [177]. Re-scrambled [177]. Realization [42]. Received [28]. Recognition [185]. Recognizing [76]. Reconstruction [192].

Record [116, 269, 302]. Recourse [250]. Recovery [65]. Rectangular [11]. Reductions [37]. Reflecting [282]. Reflection [10]. Reflexive [266]. Registration [264, 313]. Regression [117, 137, 80, 290, 212, 270, 306, 158, 241]. Regular [78, 257, 84, 196, 120, 203]. Regularity [80, 68]. Regularization [287]. Regularly [235]. Related [267, 52, 63]. Relative [300, 304]. Relaxed [74]. Reliability [182]. Remark [3]. Removal [68]. renormalizable [199]. Renormalization [38]. Renyi [244]. Representation [33, 210]. Representations [91]. Respect [314, 27]. Response [177, 299, 170]. Responses [232, 315]. Restful [114]. Restless [114]. Restricted [137, 304]. Results [208, 214, 92, 57, 268]. Ricci [179, 89, 23, 64, 4, 49, 40, 54]. Riemann [91]. Riemannian [118, 249, 161, 106]. Riesz [40]. Rigid [255, 264]. Rigidity [57, 43]. Ring [197]. Rings [131, 82, 217, 311, 83]. Risk [235]. Ritz [128]. Robust [320]. Root [93]. Root-Finding [93]. Roots [243]. Ruin [152]. Rules [216].

S [284]. S-Noetherian [284]. Samples [242, 172]. Sampling [189, 133, 126, 176, 75, 203]. Sasaki [4]. Sasakian [4]. Satisfying [284]. Scalar [23]. Scale [166, 167]. Scale-Free [166]. Scallop [296]. Scheme [312, 294]. Schrödinger [175, 205, 180, 148]. Scrambled [171, 177]. SDEs [111, 305, 293]. Second [312, 288, 46, 295, 151]. Second-Grade [151]. Second-Order [312]. Segal [38]. Segmentation [281]. Select [227]. Selection [229, 306]. Self [129, 291, 90]. Self-Exciting [291]. Self-Normalized [129, 90]. Semi [217, 236, 162]. Semi-graded [217]. Semi-stability [162]. Semi-stable [236]. Semigroups [99, 148]. Semilattices [99]. Semiparametric [165].

Sensitive [177, 171]. Sensitivity [304, 110]. Separable [107]. Sequences [216, 103, 84]. Sequential [173]. Series [210, 1, 225]. Sets [27, 222]. Several [39]. Shannon [269]. Shaped [102]. Sharing [125]. Sharp [65, 149]. Shaw [312]. Shige [308]. Shimura [140]. Short [266]. Siegel [181]. Signal [207]. Signed [311]. Similarity [37]. Simple [175, 265, 76]. Simpson [255]. Simulation [105, 133, 232]. Simultaneous [177, 37]. Single [9]. Single-Index [9]. Singular [175, 25, 228, 150, 221, 106]. Skew [73, 239, 131, 217, 283, 200, 206]. Skew-normal [283]. SLEFE [93]. SLEFE-Based [93]. Slender [24]. Slip [15]. Smale [322]. Small [140]. Sobolev [49, 40, 54]. Sofic [13]. Softwares [250]. Solids [134]. Solitons [57]. Solubility [316, 295]. Soluble [92]. Solution [180]. Solutions [10, 289, 11, 30, 324, 96, 293]. Solvability [280]. Solvable [251]. Solving [128, 259, 55, 260]. Some [208, 99, 36, 217, 18, 121, 92, 91, 124]. Source [36, 97]. Space [166, 181, 274, 221, 32]. Spaces [181, 274, 245]. Spanning [224]. Sparse [80, 306]. Spatial [254]. Spatially [277]. SPDE [16]. SPDEs [10, 147]. Special [157, 91, 168]. Spectra [161, 209]. Spectral [281, 294, 240]. Spectral-Galerkin [294]. Spectrum [45, 76, 284]. Sphere [22]. Spheres [275, 4]. Spherical [183]. Spirals [296]. Spline [169, 274, 134, 238]. Splines [216, 123, 70]. Sporadic [76]. Spreading [112]. Stability [319, 97, 162]. stable [236]. Stage [176]. State [243]. stationary [249]. Statistical [116, 135, 52, 114, 291]. Statistics [20, 269, 268]. Status [290]. Step [132]. Step-Stress [132]. Stieltjes [321]. Stochastic [137, 303, 122, 62, 288, 25, 60, 160, 283, 151]. Stokes [205, 187, 15]. Stopping [111]. Stratified [9]. Stress [132]. Strictly [109]. Strong [309]. Stronger [110]. strongly [249]. Structure [239, 248, 98, 136, 124]. Students [95]. Study [133, 220]. Sub [267, 61]. Sub-fractional [267]. Sub-linear [61]. Subgroup [98]. Subgroups [48, 66, 108, 142, 295, 138, 58, 100, 124, 204]. Submaximal [142]. Subnormal [124]. Sufficient [25]. Suitable [70]. Sum [139]. Supercloseness [11]. Supercritical [293]. Supersolvability [121]. supersolvable [66]. Supervised [276]. Support [191]. Surface [118, 256, 106]. Surfaces [89, 34, 46, 314, 67, 91, 313, 53]. Survey [189, 147, 231]. Switching [219]. Sylow [204]. Symmetric [210, 77, 51, 136]. Symmetrical [182]. Symmetricity [220]. Symplectic [46, 12, 57]. Syndetic [222]. System [74, 205, 160]. Systematic [184]. Systems [104, 322, 60, 324].

T [274, 270]. T-meshes [274]. T-Process [270]. Tables [310]. Tail [235]. Technique [250, 170, 126]. Term [135]. Test [248, 101]. Testing [81, 317, 136, 285, 203]. Tests [132]. Teugels [160]. th [317]. Their [125, 148]. Theorem [152, 282, 253, 298, 43, 278]. Theorems [36, 308]. Theory [152, 164, 86, 156, 92]. Thin [187]. Three [172, 194, 243]. Three-Component [172]. Three-Factor [194]. Three-State [243]. Threefolds [298]. Threshold [291]. Time [218, 312, 213, 111, 243, 225, 304]. Time-Accurate [312]. Time-Restricted [304]. Time-Varying [213]. Times [198]. Toeplitz [181]. Tomography [259]. Tool [296, 254]. Topological [22, 174]. Topology [275, 246]. Torelli [298]. Toric [72]. Towers [77]. trace [16]. Traces [48]. Tracking [227]. Tractrix [254]. Traffic [135]. Traffic-Congested [135]. Training [309]. Transform [226]. Transformation [233, 171]. Transformations [118]. Transforms [30, 40]. Transition [145]. Translating [57]. Transmission [232]. Traveltime [259]. Trees [224]. Triangles [286, 257]. Triebel [245]. Triply [239]. Trudinger [221, 106]. Truncated [214]. Tsallis [244]. Tutte [287]. Two [177, 271, 22, 66, 139, 42, 176, 144, 75, 159, 124, 151]. Two-Dimensional [151]. Two-Parameter [42]. Two-Phase [75]. Two-Stage [176]. Type [50, 133, 36, 236, 245, 318, 195, 188, 132, 242, 112]. Type-I [132]. Type-II [195].

Uhlenbeck [267]. Unbiased [117, 137]. Unicritical [1]. uniform [216, 20]. Unimodality [238]. Uniqueness [141, 205, 63]. Uniruled [12]. Unitary [140]. Unity [70]. Unstable [307]. Upper [181]. Upward [163]. Use [184]. Using [95, 271, 171, 93, 126, 296, 30, 114, 180, 75].

V. [44]. Value [125, 173, 130]. Valued [234]. Values [116, 302, 291]. Variable [171, 229, 314, 306]. Variables [177, 271, 152, 229, 39, 107, 268]. Variance [248, 60]. Variate [283]. Variates [126, 75]. Variation [235, 46]. Variational [128, 94, 260]. Varying [213]. Vector [322, 191]. Vectors [308]. Version [141, 110]. Vesicles [294]. Vesicular [294]. Via [309, 191, 210, 104, 229, 145]. Vine [203]. Vlasov [293]. Volatility [303]. Volume [294]. Volume-Conserved [294].

Wave [105, 24]. Weak

[157, 120, 198, 293, 106]. Weibull [188, 242, 302, 102, 154]. Weight [106]. Weighted [20, 193]. Weights [224]. Well [184, 224]. Well-Connected [224]. Well-Founded [184]. Weyl [318]. Which [124]. Whittaker [175]. Whose [251, 138, 58, 204]. Wise [107]. Within [292]. Without [250, 257]. Witt [200]. Work [44].

Yang [143, 69]. Yau [72, 5, 298]. Yudovich [50]. Yukawa [5].

Zero [311]. Zero-Divisor [311].

References

Rivera-Letelier:2013:PSU

[1] Juan Rivera-Letelier and Weixiao Shen. On Poincaré series of unicritical polynomials at the critical point. Communications in Mathematics and Statistics, 1(1):1–17, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0002-x.pdf.

Cai:2013:DFB

[2] Wei Cai, Jian Wu, and Jianguo Xin. Divergence-free H(div)-conforming hierarchical bases for magnetohydrodynamics (MHD). Communications in Mathematics and Statistics, 1(1): 19–35, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0003-9.pdf.

Ma:2013:RDE

[3] Xiaonan Ma and George Marinescu. Remark on the off-diagonal expansion of the Bergman kernel on compact Kähler manifolds. Communications in Mathematics and Statistics, 1(1):37–41, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0004-8.pdf.

Wang:2013:SRF

[4] Guofang Wang and Yongbing Zhang. The Sasaki–Ricci flow on Sasakian 3spheres. Communications in Mathematics and Statistics, 1(1):43–71, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s40304-
013-0005-7.pdf.

Sheng:2013:MFC

[5] Mao Sheng, Jinxing Xu, and Kang Zuo. Maximal families of Calabi–Yau manifolds with minimal length Yukawa coupling. Communications in Mathematics and Statistics, 1(1):73–92, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-013-0006-6.pdf.

Hsu:2013:MCE

[6] Elton P. Hsu and Karl-Theodor Sturm. Maximal coupling of Euclidean Brownian motions. Communications in Mathematics and Statistics, 1(1):93– 104, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0007-5.pdf.

Anonymous:2013:HCa

[7] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 1(1):??, March 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Tian:2013:PEK

[8] Gang Tian. Partial C^{0}-estimate for Kähler–Einstein metrics. Communications in Mathematics and Statistics, 1(2):105–113, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0011-9.pdf.

Ding:2013:LSL

[9] Kai Ding, Michael R. Kosorok, and Donglin Zeng. On the local and stratified likelihood approaches in singleindex hazards model. Communications in Mathematics and Statistics, 1(2):115–132, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0013-7.pdf.

Dalang:2013:HCS

[10] Robert C. Dalang and Tusheng Zhang. Hölder continuity of solutions of SPDEs with reflection. Communications in Mathematics and Statistics, 1(2):133–142, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0009-3.pdf.

Huang:2013:SDF

[11] Yunqing Huang and Shangyou Zhang. Supercloseness of the divergence-free finite element solutions on rectangular grids. Communications in Mathematics and Statistics, 1(2):143–162, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-013-0012-8.pdf.

Li:2013:USD

[12] Tian-Jun Li and Yongbin Ruan. Uniruled symplectic divisors. Communications in Mathematics and Statistics, 1(2):163–212, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-013-0010-x.pdf.

Kerr:2013:CIS

[13] David Kerr and Hanfeng Li. Combinatorial independence and sofic entropy. Communications in Mathematics and Statistics, 1(2):213–257, June 2013. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s40304-013-0001-y.pdf.

Anonymous:2013:HCb

[14] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 1(2):??, June 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Xiao:2013:ILN

[15] Yuelong Xiao and Zhouping Xin. On the inviscid limit of the 3D Navier–Stokes equations with generalized Navier–slip boundary conditions. Communications in Mathematics and Statistics, 1(3): 259–279, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-013-0014-6.pdf.

DaPrato:2013:FPE

[16] G. Da Prato, F. Flandoli, and M. Röckner. Fokker–Planck equations for SPDE with non-trace-class noise. Communications in Mathematics and Statistics, 1(3):281–304, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0015-5.pdf.

Chen:2013:DQM

[17] Jingyi Chen and Jiayu Li. Detecting quaternionic maps between Hyperkähler manifolds. Communications in Mathematics and Statistics, 1(3): 305–314, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0016-4.pdf.

Li:2013:CSC

[18] Jian-Shu Li and Binyong Sun. On the cohomology of some complex hyperbolic arithmetic 3-manifolds. Communications in Mathematics and Statistics, 1 (3):315–329, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0017-3.pdf.

Wang:2013:DLB

[19] Ruimin Wang, Zhouwang Yang, Ligang Liu, and Qing Chen. Discretizing Laplace–Beltrami operator from differential quantities. Communications in Mathematics and Statistics, 1(3): 331–350, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-013-0018-2.pdf.

Hu:2013:NUB

[20] Haojun Hu and Qi-Man Shao. Nonuniform Berry–Esséen bounds for weighted U-statistics and generalized L-statistics. Communications in Mathematics and Statistics, 1(3): 351–367, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-013-0019-1.pdf.

Anonymous:2013:HCc

[21] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 1(3):??, September 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Chen:2013:LET

[22] XiuXiong Chen and Meijun Zhu. Liouville energy on a topological two sphere. Communications in Mathematics and Statistics, 1(4):369–385, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s40304-013-0020-8.pdf.

Huang:2013:LNS

[23] Hongnian Huang. The L^{\frac{3}{2}} -norm of the scalar curvature under the Ricci flow on a 3-manifold. Communications in Mathematics and Statistics, 1 (4):387–392, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-013-0008-4.pdf.

Jiang:2013:LDG

[24] Jinfeng Jiang and Yan Xu. Local discontinuous Galerkin method for the impactinduced wave in a slender cylinder composed of a non-convex elastic material. Communications in Mathematics and Statistics, 1(4):393–415, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-013-0022-6.pdf.

Hafayed:2013:MFN

[25] Mokhtar Hafayed. A mean-field necessary and sufficient conditions for optimal singular stochastic control. Communications in Mathematics and Statistics, 1 (4):417–435, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0023-0.pdf.

Wu:2013:GEH

[26] Jiaxian Wu and Yi-Hu Yang. Gradient estimates and Harnack inequality for a nonlinear parabolic equation on complete manifolds. Communications in Mathematics and Statistics, 1 (4):437–464, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0026-x.pdf.

Ma:2013:CGC

[27] Xi-Nan Ma and Wei Zhang. The concavity of the Gaussian curvature of the convex level sets of p-harmonic functions with respect to the height. Communications in Mathematics and Statistics, 1 (4):465–489, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0025-y.pdf.

Anonymous:2013:ECI

[28] Anonymous. Erratum concerning incorrect received dates in eight articles. Communications in Mathematics and Statistics, 1(4):491–493, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s40304-013-0021-7.pdf.

Anonymous:2013:HCd

[29] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 1(4):??, December 2013. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Mohan:2014:SNF

[30] J. Jagan Mohan and G. V. S. R. Deekshitulu. Solutions of nabla fractional difference equations using N-transforms. Communications in Mathematics and Statistics, 2(1):1–16, March 2014. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s40304-014-0027-9.pdf.

Chen:2014:IO

[31] Mu-Fa Chen and Xu Zhang. Isospectral operators. Communications in Mathematics and Statistics, 2(1):17– 32, March 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0028-8.pdf.

Zong:2014:SCC

[32] Hong R. Zong. On the space of conics on complete intersections. Communications in Mathematics and Statistics, 2(1):33–45, March 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0029-7.pdf.

Benth:2014:RID

[33] Fred Espen Benth and Paul Krühner. Representation of infinite-dimensional forward price models in commodity markets. Communications in Mathematics and Statistics, 2(1):47–106, March 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-014-0030-1.pdf.

Feng:2014:MMC

[34] Wenyue Feng, Zhouwang Yang, and Jiansong Deng. Moving multiple Curves/ surfaces approximation of mixed point clouds. Communications in Mathematics and Statistics, 2(1):107–124, March 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0031-0.pdf.

Anonymous:2014:HCa

[35] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 2(1):??, March 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Kumar:2014:SSC

[36] Satish Kumar, Arun Choudhary, and Arvind Kumar. Some source coding theorems and 1: 1 coding based on generalized inaccuracy measure of order α and type β. Communications in Mathematics and Statistics, 2(2):125–138, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0032-z.pdf.

Li:2014:SSR

[37] Ren-Cang Li and Qiang Ye. Simultaneous similarity reductions for a pair of matrices to condensed forms. Communications in Mathematics and Statistics, 2(2):139–153, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0033-y.pdf.

Losev:2014:LOR

[38] Andrey Losev. Local observables and renormalization in Dirac–Segal approach to QFT. Communications in Mathematics and Statistics, 2(2):155–161, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0036-8.pdf.

Wang:2014:OAS

[39] Haiyan Wang and Guangbin Ren. Octonion analysis of several variables. Communications in Mathematics and Statistics, 2(2):163–185, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0034-x.pdf.

Ye:2014:SIR

[40] Rugang Ye. Sobolev inequalities, Riesz transforms, and the Ricci flow. Communications in Mathematics and Statistics, 2(2):187–209, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-014-0035-9.pdf.

Anonymous:2014:HCb

[41] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 2(2):??, June 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Jing:2014:MRT

[42] Naihuan Jing and Ming Liu. R-matrix realization of two-parameter quantum group U_{r,s}(}↕_{n}). Communications in Mathematics and Statistics, 2(3–4): 211–230, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0037-7.pdf.

Wu:2014:NOR

[43] Guoqiang Wu and Rugang Ye. A note on Obata’s rigidity theorem. Communications in Mathematics and Statistics, 2(3–4):231–252, December 2014. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-014-0038-6.pdf.

Zhang:2014:MWV

[44] Weiping Zhang. The mathematical work of V. K. Patodi. Communications in Mathematics and Statistics, 2 (3–4):253–277, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0039-5.pdf.

Chen:2014:CDS

[45] Mu-Fa Chen. Criteria for discrete spectrum of 1D operators. Communications in Mathematics and Statistics, 2 (3–4):279–309, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0041-y.pdf.

Han:2014:SVF

[46] Xiaoli Han and Jiayu Li. The second variation of the functional L of symplectic critical surfaces in Kähler surfaces. Communications in Mathematics and Statistics, 2(3–4):311– 330, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-014-0040-z.pdf.

Chen:2014:HEN

[47] Zhen-Qing Chen and Xicheng Zhang. Hölder estimates for nonlocal–diffusion equations with drifts. Communications in Mathematics and Statistics, 2 (3–4):331–348, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-015-0042-5.pdf.

Guo:2014:BFT

[48] W. Guo, Alexander N. Skiba, and X. Tang. On boundary factors and traces of subgroups of finite groups. Communications in Mathematics and Statistics, 2(3–4):349– 361, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-015-0043-4.pdf.

Ye:2014:LSI

[49] Rugang Ye. The logarithmic Sobolev inequality along the Ricci flow: The case λ_{0}(g_{0}) = 0. Communications in Mathematics and Statistics, 2(3–4): 363–368, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-015-0044-3.pdf.

Huang:2014:EBE

[50] Aimin Huang. The 2D Euler–Boussinesq equations in planar polygonal domains with Yudovich’s type data. Communications in Mathematics and Statistics, 2(3–4):369–391, December 2014. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s40304-015-0045-2.pdf.

Liu:2014:SCC

[51] Hui Liu, Yiming Long, Wei Wang, and Ping’an Zhang. Symmetric closed characteristics on symmetric compact convex hypersurfaces in R^{8}. Communications in Mathematics and Statistics, 2 (3–4):393–411, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-015-0047-0.pdf.

Princy:2014:KFR

[52] T. Princy. Krätzel function and related statistical distributions. Communications in Mathematics and Statistics, 2 (3–4):413–429, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s40304-015-0048-z.pdf.

Zhang:2014:DGM

[53] Futao Zhang, Yan Xu, and Falai Chen. Discontinuous Galerkin methods for isogeometric analysis for elliptic equations on surfaces. Communications in Mathematics and Statistics, 2(3–4): 431–461, December 2014. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s40304-015-0049-y.pdf.

Ye:2015:LSS

[54] Rugang Ye. The logarithmic Sobolev and Sobolev inequalities along the Ricci flow. Communications in Mathematics and Statistics, 3(1):1–36, March 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0046-1.

Li:2015:ADM

[55] Yusheng Li, Xinchang Xie, and Zhouwang Yang. Alternating direction method of multipliers for solving dictionary learning models. Communications in Mathematics and Statistics, 3(1):37–55, March 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0050-5.

Huang:2015:NDM

[56] Wen Huang and Xiangdong Ye. A note on double minimality. Communications in Mathematics and Statistics, 3(1):57–61, March 2015. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0051-4.

Sun:2015:RRL

[57] Jun Sun. Rigidity results on Lagrangian and symplectic translating solitons. Communications in Mathematics and Statistics, 3(1):63–68, March 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0052-3.

Zhang:2015:FPG

[58] Qinhai Zhang, Libo Zhao, Miaomiao Li, and Yiqun Shen. Finite p-groups all of whose subgroups of index p^{3} are Abelian. Communications in Mathematics and Statistics, 3(1):69–162, March 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0053-2.

Anonymous:2015:HC

[59] Anonymous. Help & contacts. Communications in Mathematics and Statistics, 3(1):??, March 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

Hafayed:2015:MFM

[60] Mokhtar Hafayed, Moufida Tabet, and Samira Boukaf. Mean-field maximum principle for optimal control of forward–backward stochastic systems with jumps and its application to mean–variance portfolio problem. Communications in Mathematics and Statistics, 3(2):163–186, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0054-1.

Zhang:2015:DIP

[61] Li-Xin Zhang. Donsker’s invariance principle under the sub-linear expectation with an application to Chung’s law of the iterated logarithm. Communications in Mathematics and Statistics, 3(2):187–214, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0055-0.

Estep:2015:PMS

[62] Donald J. Estep and Peter L. Polyakov. On a perturbation method for stochastic parabolic PDE. Communications in Mathematics and Statistics, 3(2):215– 226, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-015-
0056-z.

Sahoo:2015:UEF

[63] Pulak Sahoo. Uniqueness of entire functions related to difference polynomials. Communications in Mathematics and Statistics, 3(2):227–238, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0057-y.

Liu:2015:CGK

[64] Jiawei Liu and Yue Wang. Convergence of the generalized Kähler–Ricci flow. Communications in Mathematics and Statistics, 3(2):239–261, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0058-x.

Liu:2015:RSF

[65] Zhao Liu, Maodong Pan, Zhouwang Yang, and Jiansong Deng. Recovery of sharp features in mesh models. Communications in Mathematics and Statistics, 3(2):263–283, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0059-9.

Guo:2015:FMN

[66] Wenbin Guo and A. S. Kondrat’ev. Finite minimal non-supersolvable groups decomposable into the product of two normal supersolvable subgroups. Communications in Mathematics and Statistics, 3(2):285–290, June 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0060-3.

Mai:2015:CEH

[67] Jiehua Mai and Song Shao. Chaotic expansive homeomorphisms on closed orientable surfaces of positive genus. Communications in Mathematics and Statistics, 3(3):291–314, September 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0061-2.

Szemeredi:2015:APD

[68] Endre Szemerédi. Arithmetic progressions, different regularity lemmas and removal lemmas. Communications in Mathematics and Statistics, 3(3):315–328, September 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0062-1.

Hu:2015:HPG

[69] Sen Hu and Peng Liu. HOM-FLY polynomial from a generalized Yang–Yang function. Communications in Mathematics and Statistics, 3(3):329–352, September 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0063-0.

Zhang:2015:LIP

[70] Jingjing Zhang and Xin Li. On the linear independence and partition of unity of arbitrary degree analysis-suitable T - splines. Communications in Mathematics and Statistics, 3(3):353–364, September 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0064-z.

Draouil:2015:DDF

[71] Olfa Draouil and Bernt Øksendal. A Donsker delta functional approach to optimal insider control and applications to finance. Communications in Mathematics and Statistics, 3(3):365–421, September 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-015-0065-y. See er-

ratum [79].

Doi:2015:DCC

[72] Mamoru Doi and Naoto Yotsutani. Doubling construction of Calabi–Yau fourfolds from toric Fano fourfolds. Communications in Mathematics and Statistics, 3(3):423–447, September 2015.

CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0066-x.

Artamonov:2015:DSP

[73] V. A. Artamonov. Derivations of skew PBW — extensions. Communications in Mathematics and Statistics, 3(4):449–457, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0067-9.

Chala:2015:NRC

[74] Adel Chala. Near-relaxed control problem of fully coupled forward– backward doubly system. Communications in Mathematics and Statistics, 3(4):459–476, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0068-8.

Vishwakarma:2015:ECE

[75] Gajendra K. Vishwakarma and Manish Kumar. An efficient class of estimators for the mean of a finite population in two-phase sampling using multi-auxiliary variates. Communications in Mathematics and Statistics, 3(4):477–489, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0069-7.

Mazurov:2015:RSA

[76] V. D. Mazurov and A. R. Moghaddamfar. Recognizing by spectrum for the automorphism groups of sporadic simple groups. Communications in Mathematics and Statistics, 3(4):491–496, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0070-1.

Fang:2015:SCH

[77] Fuquan Fang, Fengchun Lei, and Jie Wu. The symmetric commutator homology of link towers and homotopy groups of 3-manifolds. Communications in Mathematics and Statistics, 3(4):497–526, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0071-0.

Makhnev:2015:ADR

[78] A. A. Makhnev and D. V. Paduchikh. On automorphisms of distance–regular graph with intersection array {18, 15, 9; 1, 1, 10}. Communications in Mathematics and Statistics, 3(4):527–534, December 2015. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0072-z.

Draouil:2015:EDD

[79] Olfa Draouil and Bernt Øksendal. Erratum to: A Donsker Delta Functional Approach to Optimal Insider Control and Applications to Finance. Communications in Mathematics and Statistics, 3(4):535–540, December 2015. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

//link.springer.com/article/10.
1007/s40304-015-0074-x;   http://
link.springer.com/content/pdf/10.

1007/s40304-015-0074-x.pdf.  See

[71].

Dobriban:2016:RPS

[80] Edgar Dobriban and Jianqing Fan. Regularity properties for sparse regression. Communications in Mathematics and Statistics, 4(1):1–19, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0078-6.

Chen:2016:TON

[81] Jiahua Chen and Pengfei Li. Testing the order of a normal mixture in mean. Communications in Mathematics and Statistics, 4(1):21–38, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0079-5.

Dhara:2016:GDA

[82] Basudeb Dhara and Nurcan Argaç. Generalized derivations acting on multilinear polynomials in prime rings and Banach algebras. Communications in Mathematics and Statistics, 4(1):39–54, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-015-
0073-y.

Tekir:2016:AQP

[83] Unsal Tekir, Suat Koç, Kursat Hakan Oral, and Kar Ping Shum. On 2absorbing quasi–primary ideals in commutative rings. Communications in Mathematics and Statistics, 4(1):55–62, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-015-
0075-9.

Ozcag:2016:DCN

[84] Emin Özçağ, Limonka Lazarova, and Biljana Jolevska-Tuneska. Defining compositions of x_{+}^{μ}, |x|^{μ}, x^{−s}, and x^{−s} ln |x| as neutrix limit of regular sequences. Communications in Mathematics and Statistics, 4(1):63–80, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0076-8.

Li:2016:LDM

[85] Changjing Li, Quanyuan Chen, and Ting Wang. ∗-Lie derivable mappings on von Neumann algebras. Communications in Mathematics and Statistics, 4(1):81–92, March 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0077-7.

Lewis:2016:LHT

[86] James D. Lewis. Lectures on Hodge theory and algebraic cycles. Communications in Mathematics and Statistics, 4(2):93–188, June 2016. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0080-z.

Artamonov:2016:EMO

[87] Viacheslav Artamonov, Oswaldo Lezama, and William Fajardo. Extended modules and Ore extensions. Communications in Mathematics and Statistics, 4(2):189–202, June 2016. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0081-y.

Deng:2016:BMP

[88] Fang Deng, Chao Zeng, and Jiansong Deng. Boundary-mapping parametrization in isogeometric analysis. Communications in Mathematics and Statistics, 4(2):203–216, June 2016. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0082-x.

Fang:2016:EMF

[89] Shouwen Fang, Liang Zhao, and Peng Zhu. Estimates and monotonicity of the first eigenvalues under the Ricci flow on closed surfaces. Communications in Mathematics and Statistics, 4(2):217–228, June 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0083-9.

Zhang:2016:SNM

[90] Li-Xin Zhang. Self-normalized moderate deviation and laws of the iterated logarithm under G-expectation. Communications in Mathematics and Statistics, 4(2):229–263, June 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-015-0084-8.

Sun:2016:SSF

[91] Ruiran Sun. Some special families of rank-2 representations of π_{1} of compact Riemann surfaces. Communications in Mathematics and Statistics, 4(2):265–279, June 2016. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-016-0085-2.

Skiba:2016:SRT

[92] Alexander N. Skiba. On some results in the theory of finite partially soluble groups. Communications in Mathematics and Statistics, 4(3): 281–309, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0088-z.

Jiang:2016:PRF

[93] Ping Jiang, Xingqiao Wu, and Zhi Liu. Polynomial root-finding using a SLEFEbased clipping method. Communications in Mathematics and Statistics, 4 (3):311–322, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0086-1.

Treanta:2016:MFV

[94] Savin Treanta. Multiobjective fractional variational problem on higherorder jet bundles. Communications in Mathematics and Statistics, 4(3): 323–340, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0087-0.

Amissah:2016:AIS

[95] Ekow Ewusi Amissah, Nana Kena Frempong, and Emmanuel DeGraft Johnson Owusu-Ansah. Accessing individual students academic performance using random effect analysis (multilevel analysis). Communications in Mathematics and Statistics, 4(3): 341–357, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0089-y.

Zhang:2016:FSN

[96] Xicheng Zhang. Fundamental solutions of nonlocal Hörmander’s operators. Communications in Mathematics and Statistics, 4(3):359–402, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0090-5.

Qin:2016:SEI

[97] Xue Qin and Shumin Li. A stability estimate for an inverse problem of determining a coefficient in a hyperbolic equation with a point source. Communications in Mathematics and Statistics, 4(3):403–421, September 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0091-4.

Qiao:2016:HDD

[98] Shouhong Qiao, Guohua Qian, and Yanming Wang. How does diagonal subgroup embedding determine the structure of a group? Communications in Mathematics and Statistics, 4(4): 423–433, December 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0092-3.

He:2016:MPC

[99] Yong He, Wei Tian, and Kar Ping Shum. P-condense and CP-condensing operators on semilattices, join-complete lattices and some inverse semigroups. Communications in Mathematics and Statistics, 4(4):435–447, December 2016. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s40304-016-0093-
                                        2.

Zhang:2016:EPS

[100] Jia Zhang, Long Miao, and Baojun Li. On m-embedded primary subgroups of finite groups. Communications in Mathematics and Statistics, 4 (4):449–458, December 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s40304-016-0094-1.

Noughabi:2016:ELR

[101] Hadi Alizadeh Noughabi. Empirical likelihood ratio-based goodness-of-fit test for the Laplace distribution. Communications in Mathematics and Statistics, 4 (4):459–471, December 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s40304-016-0095-0.

Singh:2016:APW

[102] Bhupendra Singh. An additive Perks– Weibull model with bathtub-shaped hazard rate function. Communications in Mathematics and Statistics, 4 (4):473–493, December 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0096-z.

Galmak:2016:INS

[103] Alexander M. Gal’mak and Viktoria A.

Kovaleva. On identities and m-neutral sequences of n-ary groups. Communications in Mathematics and Statistics, 4 (4):495–508, December 2016. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0097-y.

E:2017:PML

[104] Weinan E. A proposal on machine learning via dynamical systems. Communications in Mathematics and Statistics, 5(1):1–11, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-017-0103-z.

Duan:2017:NSC

[105] Yali Duan, Linghua Kong, and Min Guo. Numerical simulation of a class of nonlinear wave equations by lattice Boltzmann method. Communications in Mathematics and Statistics, 5(1):13–35, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0098-x.

Zhu:2017:WTM

[106] Xiaobao Zhu. A weak Trudinger–Moser inequality with a singular weight on a compact Riemannian surface. Communications in Mathematics and Statistics, 5(1):37–57, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-016-0099-9.

Yang:2017:ADM

[107] Jiaojiao Yang, Yusheng Li, Xinchang Xie, and Zhouwang Yang. Alternating direction method for separable variables under pair-wise constraints. Communications in Mathematics and Statistics, 5(1):59–82, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-017-0100-2.

Guo:2017:FGP

[108] Wenbin Guo, Chenchen Cao, Alexander N. Skiba, and Darya A. Sinitsa. Finite groups with H-permutable subgroups. Communications in Mathematics and Statistics, 5(1):83–92, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s40304-017-0101-1.

Huang:2017:SEM

[109] Wen Huang, Song Shao, and Xiangdong Ye. Strictly ergodic models under face and parallelepiped group actions. Communications in Mathematics and Statistics, 5(1):93–122, March 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s40304-017-0102-0.

Zou:2017:SVS

[110] Yong Zou. Stronger version sensitivity, almost finite to one extension and maximal pattern entropy. Communications in Mathematics and Statistics, 5(2):123–139, June 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0104-y.

Guerdouh:2017:FBS

[111] Dalila Guerdouh and Nabil Khelfallah. Forward-backward SDEs driven by Lévy process in stopping time duration. Communications in Mathematics and Statistics, 5(2):141–157, June 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0105-x.

Panahi:2017:EMG

[112] Hanieh Panahi. Estimation methods for the generalized inverted exponential distribution under Type II progressively hybrid censoring with application to spreading of micro-drops data. Communications in Mathematics and Statistics, 5(2):159–174, June 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0106-9.

Zhang:2017:HGF

[113] Zhe Zhang and Qin Yue. Hilbert genus fields of imaginary biquadratic fields. Communications in Mathematics and Statistics, 5(2):175–197, June 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0107-8.

Priyadarshini:2017:USM

[114] Poonam Priyadarshini and Soubhik Chakraborty. Using statistical modeling, rate of change of pitch and interonset interval to distinguish between restful and restless ragas. Communications in Mathematics and Statistics, 5(2):199–212, June 2017. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0108-7.

Joshi:2017:NEF

[115] Rajesh Joshi and Satish Kumar. A new exponential fuzzy entropy of order- (α, β) and its application in multiple attribute decision-making problems. Communications in Mathematics and Statistics, 5(2):213–229, June 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0109-6.

Kumar:2017:SIE

[116] Devendra Kumar, Tanujit Dey, and Sanku Dey. Statistical inference of exponentiated moment exponential distribution based on lower record values. Communications in Mathematics and Statistics, 5(3):231–260, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0110-0.

Arumairajan:2017:MAU

[117] Sivarajah Arumairajan and Pushpakanthie Wijekoon. Modified almost unbiased Liu estimator in linear regression model. Communications in Mathematics and Statistics, 5(3):261–276, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0111-z.

Ai:2017:FGT

[118] Wanjun Ai. The flow of gauge transformations on Riemannian surface with boundary. Communications in Mathematics and Statistics, 5(3):277–316, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0112-y.

Zhang:2017:MCP

[119] Chun Zhang and Ligang Liu. Manifold construction over polyhedral mesh. Communications in Mathematics and Statistics, 5(3):317–333, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0113-x.

Tao:2017:WIP

[120] Xinkai Tao, Boyuan Liu, and Xinmin Hou. Weak internal partition of regular graphs. Communications in Mathematics and Statistics, 5(3):335–338, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0114-9.

Miao:2017:SCS

[121] Liyun Miao and Yangming Li. Some criteria for p-supersolvability of a finite group. Communications in Mathematics and Statistics, 5(3):339–348, September 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0115-8.

E:2017:DLB

[122] Weinan E, Jiequn Han, and Arnulf Jentzen. Deep learning-based numerical methods for high-dimensional parabolic partial differential equations and backward stochastic differential equations. Communications in Mathematics and Statistics, 5(4):349–380, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0117-6.

Zhu:2017:MBP

[123] Yuanpeng Zhu and Falai Chen. Modified bases of PHT-splines. Communications in Mathematics and Statistics, 5(4):381–397, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0116-7.

Zhang:2017:SFG

[124] Xinjian Zhang and Yong Xu. The structure of a finite group which is the product of two subgroups with some subnormal subgroups. Communications in Mathematics and Statistics, 5(4):399–405, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0118-5.

Majumder:2017:MFS

[125] Sujoy Majumder. Meromorphic functions sharing one value with their derivatives concerning the difference operator. Communications in Mathematics and Statistics, 5(4):407–427, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0119-4.

Kumar:2017:EMD

[126] Manish Kumar and Gajendra K. Vishwakarma. Estimation of mean in double sampling using exponential technique on multi-auxiliary variates. Communications in Mathematics and Statistics, 5(4):429–445, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0120-y.

Shi:2017:NG

[127] Wujie Shi and Heng Lv. A note of CP_{2} groups. Communications in Mathematics and Statistics, 5(4):447–451, December 2017. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0121-x.

E:2018:DRM

[128] Weinan E and Bing Yu. The deep Ritz method: A deep learning-based numerical algorithm for solving variational problems. Communications in Mathematics and Statistics, 6(1):1–12, March 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0127-z.

Fan:2018:BEB

[129] Xiequan Fan and Qi-Man Shao. Berry– Esseen bounds for self-normalized martingales. Communications in Mathematics and Statistics, 6(1):13–27, March 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0122-9.

Zhang:2018:QBV

[130] Tao Zhang, Changliang Zhou, and Chunqin Zhou. Quantization of the blow-up value for the Liouville equation with exponential Neumann boundary condition. Communications in Mathematics and Statistics, 6(1):29–48, March 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-017-0126-5.

DeFilippis:2018:CGS

[131] Vincenzo De Filippis and Feng Wei. Centralizers of X-generalized skew derivations on multilinear polynomials in prime rings. Communications in Mathematics and Statistics, 6(1):49–71, March 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s40304-017-
                                        0125-6.

Chandra:2018:AOP

[132] N. Chandra and Mashroor Ahmad Khan. Analysis and optimum plan for 3-step step-stress accelerated life tests with Lomax model under progressive Type-I censoring. Communications in Mathematics and Statistics, 6(1):73–90, March 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-017-0123-8.

Khan:2018:PER

[133] Hina Khan, Muhammad Aslam, and Masood Amjad Khan. Properties of exponential ratio type estimators in equal probability sampling: A simulation study. Communications in Mathematics and Statistics, 6(1):91–118, March 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-017-0124-7.

Song:2018:CAI

[134] Yanzhi Song, Yixin Luo, Yuan Liu, Jiansong Deng, and Zhouwang Yang. Compression algorithm for implicit 3D B-spline solids. Communications in Mathematics and Statistics, 6(2):119– 140, June 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-018-
0128-y.

Narayan:2018:LTS

[135] Tripta Narayan, Tanushree Bhattacharya, Soubhik Chakraborty, and Swapan Konar. Long-term statistical characteristics of air pollutants in a traffic-congested area of Ranchi, India. Communications in Mathematics and Statistics, 6(2):141–162, June 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0129-x.

Tsukada:2018:HTI

[136] Shin ichi Tsukada. Hypothesis testing for independence under blocked compound symmetric covariance structure. Communications in Mathematics and Statistics, 6(2):163–184, June 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0130-4.

Arumairajan:2018:SRM

[137] S. Arumairajan. On the stochastic restricted modified almost unbiased Liu estimator in linear regression model. Communications in Mathematics and Statistics, 6(2):185–206, June 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0131-3.

Wang:2018:FGW

[138] Lifang Wang. Finite 2-groups whose number of subgroups of each order are at most 2^{4}. Communications in Mathematics and Statistics, 6(2):207–226, June 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0133-1.

Hu:2018:HHH

[139] Ze-Chun Hu and Wei Sun. Hunt’s hypothesis (H) for the sum of two independent Lévy processes. Communications in Mathematics and Statistics, 6(2):227–247, June 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0136-y.

Chen:2018:OCS

[140] Ke Chen, Xin Lu, and Kang Zuo. The Oort conjecture for Shimura curves of small unitary rank. Communications in Mathematics and Statistics, 6(3):249–268, September 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s40304-018-0155-8;   http://

link.springer.com/content/pdf/10.
1007/s40304-018-0155-8.pdf.

Gonzalez:2018:EUF

[141] María del Mar González, YanYan Li, and Luc Nguyen. Existence and uniqueness to a fully nonlinear version of the Loewner–Nirenberg problem. Communications in Mathematics and Statistics, 6(3):269–288, September 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

//link.springer.com/article/10.
1007/s40304-018-0150-0;   http://
link.springer.com/content/pdf/10.
1007/s40304-018-0150-0.pdf.

Guo:2018:PSM

[142] Wenbin Guo and Danila O. Revin.

Pronormality and submaximal X - subgroups on finite groups. Communications in Mathematics and Statistics, 6(3):289–317, September 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

//link.springer.com/article/10.
1007/s40304-018-0154-9;   http://
link.springer.com/content/pdf/10.
1007/s40304-018-0154-9.pdf.

Li:2018:NCE

[143] Jiayu Li, Chuanjing Zhang, and Xi Zhang. A note on curvature estimate of the Hermitian–Yang–Mills flow. Communications in Mathematics and Statistics, 6(3):319–358, September 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

//link.springer.com/article/10.
1007/s40304-018-0135-z;   http://
link.springer.com/content/pdf/10.
1007/s40304-018-0135-z.pdf.

Ma:2018:LDG

[144] Tian Ma and Yan Xu. Local discontinuous Galerkin methods for the two–dimensional Camassa–Holm equation. Communications in Mathematics and Statistics, 6(3):359–388, September 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s40304-018-
                                        0140-2;   http://link.springer.
                                        com/content/pdf/10.1007/s40304-
                                        018-0140-2.pdf.

Wang:2018:PTK

[145] Jun Wang and Jian Zhou. Phase transition of Kähler–Einstein metrics via moment maps. Communications in Mathematics and Statistics, 6(3):389–415, September 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s40304-018-0153-x;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s40304-018-0153-x.pdf.

Funaki:2018:HLE

[146] Tadahisa Funaki. Hydrodynamic limit for exclusion processes. Communications in Mathematics and Statistics, 6(4):417–480, December 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0161-x.

Song:2018:SCG

[147] Jian Song. SPDEs with colored Gaussian noise: A survey. Communications in Mathematics and Statistics, 6(4):481–492, December 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0160-y.

Wang:2018:DHK

[148] Jian Wang. On-diagonal heat kernel estimates for Schrödinger semigroups and their application. Communications in Mathematics and Statistics, 6(4):493–508, December 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0163-8.

Xu:2018:SCN

[149] Weijun Xu. Sharp convergence of nonlinear functionals of a class of Gaussian random fields. Communications in Mathematics and Statistics, 6(4):509–532, December 2018. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

//link.springer.com/article/10.
1007/s40304-018-0162-9;   http://
link.springer.com/content/pdf/10.
1007/s40304-018-0162-9.pdf.

Zhang:2018:SBD

[150] Xicheng Zhang and Guohuan Zhao.

Singular Brownian diffusion processes. Communications in Mathematics and Statistics, 6(4):533–581, December 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0164-7.

Zheng:2018:MDS

[151] Wuting Zheng, Jianliang Zhai, and Tusheng Zhang. Moderate deviations for stochastic models of two-dimensional second-grade fluids driven by Lévy noise. Communications in Mathematics and Statistics, 6(4):583–612, December 2018. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0165-6.

Chen:2019:EBT

[152] Yu Chen, Dan Chen, and Wenxue Gao. Extensions of Breiman’s theorem of product of dependent random variables with applications to ruin theory. Communications in Mathematics and Statistics, 7(1):1–23, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0132-2.

Tang:2019:NCQ

[153] Juping Tang, Jia Zhang, and Long Miao. New criteria for quasi-F-groups. Communications in Mathematics and Statistics, 7(1):25–32, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0134-0.

Ye:2019:GWC

[154] Gen Ye and Songjian Wang. The Gamma/Weibull customer lifetime model. Communications in Mathematics and Statistics, 7(1):33–59, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0137-x.

Liu:2019:BQF

[155] Yang Liu and Yi Ouyang. On binary quadratic forms modulo n. Communications in Mathematics and Statistics, 7(1):61–67, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0141-1.

Makhnev:2019:IPG

[156] A. A. Makhnev, M. P. Golubyatnikov, and Wenbin Guo. Inverse problems in graph theory: Nets. Communications in Mathematics and Statistics, 7(1):69–83, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0159-4.

Huang:2019:SFW

[157] Wen Huang and Leiye Xu. Special flow, weak mixing and complexity. Communications in Mathematics and Statistics, 7(1):85–122, March 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0166-5.

Yang:2019:BNQ

[158] Hang Yang, Zhuojian Chen, and Weiping Zhang. Bayesian nonlinear quantile regression approach for longitudinal ordinal data. Communications in Mathematics and Statistics, 7(2):123– 140, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-018-
0148-7.

Vovan:2019:CTP

[159] Tai Vovan, Loc Tranphuoc, and Ha Chengoc. Classifying two populations by Bayesian method and applications. Communications in Mathematics and Statistics, 7(2):141–161, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0139-8.

Muthukumar:2019:MFI

[160] P. Muthukumar and R. Deepa. Meanfield, infinite horizon, optimal control of nonlinear stochastic delay system governed by Teugels martingales associated with Lévy processes. Communications in Mathematics and Statistics, 7(2):163–180, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0143-z.

Qin:2019:NIR

[161] Yuguo Qin. Non-isometric Riemannian G-manifolds with equal equivariant spectra. Communications in Mathematics and Statistics, 7(2):181–190, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0149-6.

Wang:2019:SSH

[162] Ruixin Wang and Chuanjing Zhang. Semi-stability for holomorphic pairs on compact Gauduchon manifolds. Communications in Mathematics and Statistics, 7(2):191–206, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0152-y.

Lu:2019:CCU

[163] Xuexing Lu and Yu Ye. Combinatorial characterization of upward planarity. Communications in Mathematics and Statistics, 7(2):207–223, June 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0169-2.

Jost:2019:GDM

[164] Jürgen Jost and Sylvia Yaptieu. A generalized discrete Morse–Floer theory. Communications in Mathematics and Statistics, 7(3):225–252, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-018-
0167-4;   http://link.springer.
com/content/pdf/10.1007/s40304-
018-0167-4.pdf.

Liu:2019:BJS

[165] Meimei Liu, Weiping Zhang, and Yu Chen. Bayesian joint semiparametric mean-covariance modeling for longitudinal data. Communications in Mathematics and Statistics, 7(3):253–267, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0138-9.

Deprez:2019:SFP

[166] Philippe Deprez and Mario V. Wüthrich.

Scale-free percolation in continuum space. Communications in Mathematics and Statistics, 7(3):269–308, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0142-0.

Tripathi:2019:EES

[167] Yogesh Mani Tripathi, Constantinos Petropoulos, and Farha Sultana. Estimating an exponential scale parameter under double censoring. Communications in Mathematics and Statistics, 7(3):309–328, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0144-y.

Wang:2019:NPS

[168] Jun Wang. The Neumann problem for special Lagrangian equations with critical phase. Communications in Mathematics and Statistics, 7(3):329–361, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0157-6.

Kang:2019:NMD

[169] Hongmei Kang and Xin Li. A new method to design cubic Pythagoreanhodograph spline curves with control polygon. Communications in Mathematics and Statistics, 7(3):363–381, September 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0158-5.

Ewemooje:2019:AED

[170] O. S. Ewemooje, F. B. Adebola, and G. N. Amahia. Alternative estimator in dichotomous randomized response technique. Communications in Mathematics and Statistics, 7(4):383–400, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0145-x.

Grover:2019:ESE

[171] Lovleen Kumar Grover and Amanpreet Kaur. An efficient scrambled estimator of population mean of quantitative sensitive variable using general linear transformation of non-sensitive auxiliary variable. Communications in Mathematics and Statistics, 7(4):401–415, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0146-9.

Feroze:2019:TCM

[172] Navid Feroze and Muhammad Aslam.

Three-component mixture of Rayleigh model under doubly censored samples: a Bayesian look. Communications in Mathematics and Statistics, 7(4):417–443, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0147-8.

Yang:2019:SBD

[173] Yunli Yang, Jing Kong, Lu Yang, and Zhouwang Yang. Sequential big data-based macroeconomic forecast for industrial value added. Communications in Mathematics and Statistics, 7(4):445–457, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00177-4.

Yang:2019:NII

[174] Tianhui Yang, Ammar Qarariyah, Hongmei Kang, and Jiansong Deng. Numerical integration over implicitly defined domains with topological guarantee. Communications in Mathematics and Statistics, 7(4):459–474, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00178-3.

Cai:2019:SSW

[175] Yan an Cai and Xiufu Zhang. Simple singular Whittaker modules over the Schrödinger algebra. Communications in Mathematics and Statistics, 7(4):475–483, December 2019. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00180-9.

Latpate:2020:TSN

[176] R. V. Latpate and J. K. Kshirsagar. Two-stage negative adaptive cluster sampling. Communications in Mathematics and Statistics, 8(1):1–21, March 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0151-z.

Ahmed:2020:FRS

[177] Segun Ahmed, Stephen A. Sedory, and Sarjinder Singh. Forcibly rescrambled randomized response model for simultaneous estimation of means of two sensitive variables. Communications in Mathematics and Statistics, 8(1):23–45, March 2020. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0156-7.

Li:2020:PFA

[178] Yaguang Li and Baisuo Jin. Pairwise fusion approach incorporating prior constraint information. Communications in Mathematics and Statistics, 8(1):47–62, March 2020. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-0168-3.

Chen:2020:ERF

[179] Youmin Chen. On expansions of Ricci flat ALE metrics in harmonic coordinates about the infinity. Communications in Mathematics and Statistics, 8(1):63–90, March 2020. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00183-6.

Qarariyah:2020:NSS

[180] Ammar Qarariyah, Fang Deng, Tianhui Yang, and Jiansong Deng. Numerical solution for Schrödinger eigenvalue problem using isogeometric analysis on implicit domains. Communications in Mathematics and Statistics, 8(1):91–111, March 2020. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00186-3.

Liu:2020:PTO

[181] Congwen Liu and Jiajia Si. Positive Toeplitz operators on the Bergman spaces of the Siegel upper half–space. Communications in Mathematics and Statistics, 8(1):113–134, March 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-019-00187-2.

Prasad:2020:RAS

[182] Penti Hari Prasad, T. Sumathi Uma Maheswari, and J. Shirisha. Reliability analysis of symmetrical columns with eccentric loading from Lindley distribution. Communications in Mathematics and Statistics, 8(2):135–149, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-0170-9.

Moghimbeygi:2020:SLD

[183] M. Moghimbeygi and M. Golalizadeh. Spherical logistic distribution. Communications in Mathematics and Statistics, 8(2):151–166, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-00171-2.

DaSilva:2020:BBS

[184] Stéphane Blondeau Da Silva. Benford or not Benford: A systematic but not always well-founded use of an elegant law in experimental fields. Communications in Mathematics and Statistics, 8(2):167–201, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-018-00172-1.

UlRahman:2020:APD

[185] Jamshaid Ul Rahman, Qing Chen, and Zhouwang Yang. Additive parameter for deep face recognition. Communications in Mathematics and Statistics, 8(2):203–217, June 2020.

CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00198-z.

Shen:2020:CMH

[186] Zhenghan Shen and Pan Zhang. Canonical metrics on holomorphic filtrations over compact Hermitian manifolds. Communications in Mathematics and Statistics, 8(2):219–237, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00199-y.

Sun:2020:IIN

[187] Yongzhong Sun and Shifang Wang. Inhomogeneous incompressible Navier– Stokes equations on thin domains. Communications in Mathematics and Statistics, 8(2):239–253, June 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00202-6.

Bdair:2020:EPF

[188] O. M. Bdair, R. R. Abu Awwad, G. K.

Abufoudeh, and M. F. M. Naser. Estimation and prediction for flexible Weibull distribution based on progressive Type II censored data. Communications in Mathematics and Statistics, 8(3):255–277, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-00173-0.

Enang:2020:ECC

[189] Ekaette I. Enang and Etebong P.

Clement. An efficient class of calibration ratio estimators of domain mean in survey sampling. Communications in Mathematics and Statistics, 8(3):279–293, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-00174-z.

Alizadeh:2020:OLL

[190] Morad Alizadeh, Lazhar Benkhelifa, Mahdi Rasekhi, and Bistoon Hosseini. The odd log–logistic generalized Gompertz distribution: Properties, applications and different methods of estimation. Communications in Mathematics and Statistics, 8(3):295–317, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-00175-y.

Zhou:2020:MCF

[191] Xuan Zhou, Yuanjia Wang, and Donglin Zeng. Multicategory classification via forward–backward support vector machine. Communications in Mathematics and Statistics, 8(3):319–339, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00179-2.

Yuan:2020:DIR

[192] Zijian Yuan, Yanzhi Song, Falai Chen, and Zhouwang Yang. Diamond inclusion reconstruction. Communications in Mathematics and Statistics, 8(3):341–359, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00206-2.

Uwitije:2020:NWA

[193] Rongin Uwitije, Xuhui Wang, Ammar Qarariyah, and Jiansong Deng. Nonlinear weighted average and blossoming. Communications in Mathematics and Statistics, 8(3):361–378, September 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s40304-020-
00208-5;   http://link.springer.
com/content/pdf/10.1007/s40304-
020-00208-5.pdf.

Goard:2020:CFF

[194] Joanna Goard. Closed-form formulae for European options under three-factor models. Communications in Mathematics and Statistics, 8(4):379–408, December 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-018-00176-x.

Alma:2020:ECE

[195] Özlem Gürünlü Alma and Reza Arabi Belaghi. Estimation in the complementary exponential geometric distribution based on progressive Type-II censored data. Communications in Mathematics and Statistics, 8(4):409–441, December 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00181-8.

Saikia:2020:IFC

[196] Nipen Saikia. Infinite families of congruences for 3-regular partitions with distinct odd parts. Communications in Mathematics and Statistics, 8(4):443–451, December 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-019-00182-7.

Suo:2020:QPH

[197] Jingjing Suo. A quasiconcavity property for the heat equation in a convex ring. Communications in Mathematics and Statistics, 8(4):453–462, December 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-020-00207-6.

Xu:2020:WHB

[198] Leiye Xu and Junren Zheng. Weak horseshoe with bounded-gap-hitting times. Communications in Mathematics and Statistics, 8(4):463–472, December 2020. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-020-00209-4.

Ji:2020:CFL

[199] Haoyang Ji and Simin Li. On the combinatorics of Fibonacci-like nonrenormalizable maps. Communications in Mathematics and Statistics, 8(4):473–496, December 2020. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s40304-020-00210-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s40304-020-00210-x.pdf.

Reyes:2021:MCS

[200] Armando Reyes and Camilo Rodríguez. The McCoy condition on skew Poincaré– Birkhoff–Witt extensions. Communications in Mathematics and Statistics, 9(1):1–21, March 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00184-5.

Tarvirdizade:2021:NEC

[201] Bahman Tarvirdizade and Mohammad Ahmadpour. A new extension of Chen distribution with applications to lifetime data. Communications in Mathematics and Statistics, 9(1):23–38, March 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00185-4.

Mostafanasab:2021:CDC

[202] Hojjat Mostafanasab and Ahmad Yousefian Darani. On cyclic DNA codes over F_{2} + uF_{2} + u^{2}F_{2}. Communications in Mathematics and Statistics, 9(1):39–52, March 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00188-1.

Yuan:2021:PPP

[203] Zhenfei Yuan and Taizhong Hu. pyvine:

The Python package for regular vine copula modeling, sampling and testing. Communications in Mathematics and Statistics, 9(1):53–86, March 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

http://link.springer.com/article/
10.1007/s40304-019-00195-2.

Zheng:2021:NPF

[204] Tao Zheng and Xiuyun Guo. The normalizer property for finite groups whose Sylow 2-subgroups are Abelian. Communications in Mathematics and Statistics, 9(1):87–99, March 2021.

CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-020-00211-w.

Huang:2021:LEU

[205] Jiaxi Huang. Local existence and uniqueness of Navier–Stokes–Schrödinger system. Communications in Mathematics and Statistics, 9(1):101–118, March 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s40304-020-00214-7.

Reyes:2021:RKC

[206] Armando Reyes and Héctor Suárez. Radicals and Köthe’s conjecture for skew PBW extensions. Communications in Mathematics and Statistics, 9(2):119– 138, June 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s40304-019-
                                        00189-0.

Ding:2021:NEE

[207] Shu Ding, Yuehua Wu, and Kwok-Wai Tam. Notes on M-estimation in exponential signal models. Communications in Mathematics and Statistics, 9(2):139–151, June 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-019-00190-7.

Das:2021:SOR

[208] Sangita Das and Suchandan Kayal. Some ordering results for the Marshall and Olkin’s family of distributions. Communications in Mathematics and Statistics, 9(2):153–179, June 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00191-
6.

Tsonev:2021:SFG

[209] D. M. Tsonev and R. R. Mesquita.

On the spectra of a family of geometric operators evolving with geometric flows. Communications in Mathematics and Statistics, 9(2):181–202, June 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00215-
6.

Ding:2021:SRJ

[210] Yujia Ding and Qidi Peng. Series representation of jointly SαS distribution via symmetric covariations. Communications in Mathematics and Statistics, 9(2):203–238, June 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00216-
5.

Fang:2021:CFM

[211] Xingui Fang and Lijian An. A classification of finite metahamiltonian pgroups. Communications in Mathematics and Statistics, 9(2):239–260, June 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00229-
0.

Wang:2021:BEE

[212] Zhanfeng Wang and Kai Li. Bayesian estimation for the extended t-process regression models with independent errors. Communications in Mathematics and Statistics, 9(3):261–272, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00192-
5.

Ghezal:2021:QPT

[213] Ahmed Ghezal. QMLE for periodic time-varying asymmetric log GARCH models. Communications in Mathematics and Statistics, 9(3):273–297, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00193-
4.

Moran-Vasquez:2021:NRT

[214] Raúl Alejandro Morán-Vásquez and Silvia L. P. Ferrari. New results on truncated elliptical distributions. Communications in Mathematics and Statistics, 9(3):299–313, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00194-
3.

Wu:2021:MLC

[215] Xiaosheng Wu and Pengzhen Yang. The most likely common difference of arithmetic progressions among primes. Communications in Mathematics and Statistics, 9(3):315–329, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00218-
3.

Chen:2021:EGQ

[216] Peng Chen and Xin Li. Explicit Gaussian quadrature rules for C^{1} cubic splines with non-uniform knot sequences. Communications in Mathematics and Statistics, 9(3):331–345, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00220-
9.

Lezama:2021:SOP

[217] Oswaldo Lezama. Some open problems in the context of skew PBW extensions and semi-graded rings. Communications in Mathematics and Statistics, 9(3):347–378, September 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00238-
7.

Bibi:2021:FDE

[218] Abdelouahab Bibi and Fateh Merahi. Frequency-domain estimation of continuous-time bilinear processes. Communications in Mathematics and Statistics, 9(4):379–403, December 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00196-
1.

Bibi:2021:QEA

[219] Abdelouahab Bibi. QML estimation of asymmetric Markov switching GARCH(p, q) processes. Communications in Mathematics and Statistics, 9(4):405–438, December 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-019-00197-0.

Toulias:2021:IDG

[220] Thomas L. Toulias and Christos P. Kitsos. Information divergence and the generalized normal distribution: a study on symmetricity. Communications in Mathematics and Statistics, 9(4):439–465, December 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-019-00200-8.

Zhou:2021:SMT

[221] Changliang Zhou and Chunqin Zhou. Singular Moser–Trudinger inequality involving L^{n} norm in the entire Euclidean space. Communications in Mathematics and Statistics, 9(4):467–501, December 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-020-00227-
                                        2.

Qiu:2021:MMC

[222] Jiahao Qiu and Jianjie Zhao. Multidimensional multiplicative combinatorial properties of dynamical syndetic sets. Communications in Mathematics and Statistics, 9(4):503–519, December 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-020-00230-
                                        7.

Li:2021:HAF

[223] Baojun Li and Lü Gong. On fhypercentral actions of finite group. Communications in Mathematics and Statistics, 9(4):521–533, December 2021. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00232-
5.

Ganesan:2022:MST

[224] Ghurumuruhan Ganesan. Minimum spanning trees across well-connected cities and with location-dependent weights. Communications in Mathematics and Statistics, 10(1):1–50, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00201-
7.

Vovan:2022:FTS

[225] Tai Vovan and Thuy Lethithu. A fuzzy time series model based on improved fuzzy function and cluster analysis problem. Communications in Mathematics and Statistics, 10(1):51–66, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00203-
5.

Prasad:2022:PCP

[226] Akhilesh Prasad and Manoj Kumar Singh. Product and commutators of pseudo-differential operators involving Fourier–Jacobi transform. Communications in Mathematics and Statistics, 10(1):67–84, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-019-00204-
4.

Torkaman:2022:TIS

[227] Parisa Torkaman. Tracking interval to select an optimal model among non-nested copula functions. Communications in Mathematics and Statistics, 10(1):85–99, March 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-019-00205-3.

Yang:2022:EEA

[228] Saisai Yang, Chen Wang, and Tusheng Zhang. Elliptic equations associated with Brownian motion with singular drift. Communications in Mathematics and Statistics, 10(1):101–122, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-020-00213-
                                        8.

Huang:2022:VSH

[229] Xudong Huang, Nana Bao, and Guanpeng Wang. Variable selection in highdimensional error-in-variables models via controlling the false discovery proportion. Communications in Mathematics and Statistics, 10(1):123–151, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-020-00233-
                                        4.

Zhang:2022:GFE

[230] Chi Zhang, Wenbin Guo, and A-Ming Liu. On a generalisation of finite T - groups. Communications in Mathematics and Statistics, 10(1):153–162, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00240-
z.

Wang:2022:SDL

[231] He Wang and Juyong Zhang. A survey of deep learning-based mesh processing. Communications in Mathematics and Statistics, 10(1):163–194, March 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00246-
7.

Kumar:2022:ENF

[232] Abhishek Kumar and Nilam. Effects of nonmonotonic functional responses on a disease transmission model: Modeling and simulation. Communications in Mathematics and Statistics, 10(2):195– 214, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-020-
00217-4.

Gorji:2022:LGT

[233] Hossein Gorji. Logarithmic gradient transformation and chaos expansion of Itô processes. Communications in Mathematics and Statistics, 10(2):215–231, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00219-
2.

Xiong:2022:MDP

[234] Lanyu Xiong and Fukang Zhu. Minimum density power divergence estimator for negative binomial integer-valued GARCH models. Communications in Mathematics and Statistics, 10(2):233– 261, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-020-
00221-8.

Chen:2022:TDR

[235] Yu Chen, Jiayi Wang, and Weiping Zhang. Tail distortion risk measure for portfolio with multivariate regularly variation. Communications in Mathematics and Statistics, 10(2):263–285, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00223-
6.

Pan:2022:NBT

[236] Changpeng Pan. A note on Bogomolovtype inequality for semi-stable parabolic Higgs bundles. Communications in Mathematics and Statistics, 10(2):287– 298, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-020-
00231-6.

Lyu:2022:NCC

[237] Weiguo Lyu and Guodong Zhou. The normalized cochain complex of a nonsymmetric cyclic operad with multiplication is a Quesney homotopy BV algebra. Communications in Mathematics and Statistics, 10(2):299–330, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00234-
3.

Yong:2022:UIB

[238] Zhiguo Yong, Hongmei Kang, and Yi Gu. The unimodality of initial B-spline approximations in spline fitting. Communications in Mathematics and Statistics, 10(2):331–352, June 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00235-
2.

Ballester-Bolinches:2022:TFG

[239] A. Ballester-Bolinches and R. Esteban-Romero. Triply factorised groups and the structure of skew left braces. Communications in Mathematics and Statistics, 10(2):353–370, June 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00239-6.

Jost:2022:SGL

[240] Jürgen Jost, Raffaella Mulas, and Florentin Münch. Spectral gap of the largest eigenvalue of the normalized graph Laplacian. Communications in Mathematics and Statistics, 10(3):371–381, September 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-020-00222-7.

Zhang:2022:REL

[241] Weiping Zhang, Dazhi Zhao, and Yu Chen. Regression estimation for longitudinal data with nonignorable intermittent nonresponse and dropout. Communications in Mathematics and Statistics, 10(3):383–411, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00224-
5.

El-Morshedy:2022:EGI

[242] M. El-Morshedy, M. S. Eliwa, A. El-Gohary, Ehab M. Almetwally, and R. El-Desokey. Exponentiated generalized inverse flexible Weibull distribution: Bayesian and non–Bayesian estimation under complete and Type II censored samples with applications. Communications in Mathematics and Statistics, 10(3):413–434, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00225-
4.

Guerry:2022:MRE

[243] Marie-Anne Guerry. Matrix roots and embedding conditions for threestate discrete-time Markov chains with complex eigenvalues. Communications in Mathematics and Statistics, 10(3):435–450, September 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-020-00226-3.

Kadian:2022:JRT

[244] Ratika Kadian and Satish Kumar. Jensen–Renyi’s–Tsallis fuzzy divergence information measure with its applications. Communications in Mathematics and Statistics, 10(3):451–482, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/

article/10.1007/s40304-020-00228-
1.

Wang:2022:DCB

[245] Fan Wang, Ziyi He, Dachun Yang, and Wen Yuan. Difference characterization of Besov and Triebel–Lizorkin spaces on spaces of homogeneous type. Communications in Mathematics and Statistics, 10(3):483–542, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00243-
w.

Zheng:2022:ITO

[246] Taining Zheng and Xin Li. Isogeometric topology optimization based on deep learning. Communications in Mathematics and Statistics, 10(3):543–564, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-021-
00253-8.

Wang:2022:GSP

[247] Zhigang Wang, Jin Guo, Inna N.

Safonova, and Alexander N. Skiba. A generalization of σ-permutability. Communications in Mathematics and Statistics, 10(3):565–579, September 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00309-
3.

Cao:2022:HDT

[248] Mingxiang Cao, Yanling Zhao, Kai Xu, Daojiang He, and Xudong Huang. A high-dimensional test for multivariate analysis of variance under a low-dimensional factor structure. Communications in Mathematics and Statistics, 10(4):581–597, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-020-00236-
1.

Amoussou:2022:NES

[249] Amour T. Gbaguidi Amoussou, Freedath Djibril Moussa, Carlos Ogouyandjou, and Mamadou Abdoul Diop. Nonparametric estimation for stationary and strongly mixing processes on Riemannian manifolds. Communications in Mathematics and Statistics, 10(4):599–621, December 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-020-00237-0.

Elsawah:2022:DOL

[250] A. M. Elsawah. Designing optimal large four-level experiments: a new technique without recourse to optimization softwares. Communications in Mathematics and Statistics, 10(4):623–652, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00241-
y.

Kondratev:2022:FNS

[251] A. S. Kondrat’ev and N. A. Minigulov. On finite non-solvable groups whose Gruenberg–Kegel graphs are isomorphic to the paw. Communications in Mathematics and Statistics, 10(4):653–667, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.

com/article/10.1007/s40304-021-
00242-x.

Liu:2022:EDO

[252] Xin Liu, Rong-Xian Yue, and Kashinath Chatterjee. D-optimal designs for hierarchical linear models with heteroscedastic errors. Communications in Mathematics and Statistics, 10 (4):669–679, December 2022. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00244-9.

Lian:2022:LTM

[253] Zeng Lian and Jianhua Zhang. Livsic theorem for matrix cocycles over an Axiom A flow. Communications in Mathematics and Statistics, 10(4):681–704, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-021-
00250-x.

Zou:2022:TOO

[254] Xiang Zou, Ke Shi, Hai-Yin Xu, and Hon yuen Tam. Tool orientation optimization based on spatial tractrix method for five-axis CNC machining with ball end cutters. Communications in Mathematics and Statistics, 10(4):705–737, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00255-
6.

Song:2022:RAE

[255] Ziyan Song. Rigid analytic p-adic Simpson correspondence for line bundles. Communications in Mathematics and Statistics, 10(4):739–756, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00256-
                                        5.

Cox:2022:BPL

[256] David A. Cox, Sonia Pérez-Díaz, and J. Rafael Sendra. On the base point locus of surface parametrizations: Formulas and consequences. Communications in Mathematics and Statistics, 10(4):757–783, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00257-
                                        4.

Makhnev:2022:DRG

[257] A. A. Makhnev, Wenbin Guo, and K. S. Efimov. Distance-regular graphs of diameter 3 without triangles with c_{2} = 2. Communications in Mathematics and Statistics, 10(4):785–792, December 2022. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00281-
                                        4.

Anonymous:2023:P

[258] Anonymous. Preface. Communications in Mathematics and Statistics, 11(1):1– 2, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s40304-023-
                                        00347-5.

Fan:2023:STT

[259] Yuwei Fan and Lexing Ying. Solving traveltime tomography with deep learning. Communications in Mathematics and Statistics, 11(1):3–19, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00329-
z.

Wang:2023:NNB

[260] Min Wang and Jianfeng Lu. Neural network-based variational methods for solving quadratic porous medium equations in high dimensions. Communications in Mathematics and Statistics, 11(1):21–57, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-023-00339-
5.

Li:2023:LIP

[261] Zhengyi Li, Bin Dong, and Yanli Wang.

Learning invariance preserving moment closure model for Boltzmann–BGK equation. Communications in Mathematics and Statistics, 11(1):59–101, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00331-
5.

Lin:2023:BGL

[262] Guochang Lin, Fukai Chen, Pipi Hu, Xiang Chen, Junqing Chen, Jun Wang, and Zuoqiang Shi. BI-GreenNet: Learning Green’s functions by boundary integral network. Communications in Mathematics and Statistics, 11(1):103–129, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/

                                        article/10.1007/s40304-023-00338-
                                        6.

Zhan:2023:BCI

[263] Zheng Zhan, Ye Zheng, Wenping Wang, and Falai Chen. Boundary correspondence for iso-geometric analysis based on deep learning. Communications in Mathematics and Statistics, 11(1):131– 150, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s40304-023-
                                        00337-7.

Feng:2023:DDG

[264] Wanquan Feng, Hongrui Cai, Junhui Hou, Bailin Deng, and Juyong Zhang. Differentiable deformation graph-based neural non-rigid registration. Communications in Mathematics and Statistics, 11(1):151–167, March 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-023-00341-
                                        x.

Grechkoseeva:2023:FGI

[265] Maria A. Grechkoseeva, Victor D. Mazurov, Wujie Shi, Andrey V. Vasil’ev, and Nanying Yang. Finite groups isospectral to simple groups. Communications in Mathematics and Statistics, 11(2):169–194, June 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-022-00288-5.

Ringel:2023:SLA

[266] Claus Michael Ringel. The short local algebras of dimension 6 with nonprojective reflexive modules. Communications in Mathematics and Statistics, 11(2):195–227, June 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-023-00343-9.

Cai:2023:MSF

[267] Chunhao Cai, Qinghua Wang, and Weilin Xiao. Mixed sub-fractional Brownian motion and drift estimation of related Ornstein–Uhlenbeck process. Communications in Mathematics and Statistics, 11(2):229–255, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00245-
8.

Zhang:2023:ORL

[268] Yiying Zhang, Yanni Hu, and Peng Zhao. Ordering results on largest order statistics from multiple-outlier gamma variables. Communications in Mathematics and Statistics, 11(2):257–282, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00247-
6.

Kumar:2023:QBS

[269] Vikas Kumar and Bhawna Dangi.

Quantile-based Shannon entropy for record statistics. Communications in Mathematics and Statistics, 11(2):283– 306, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-021-
00248-5.

Wang:2023:CPR

[270] Zhanfeng Wang, Yuewen Lv, and Yaohua Wu. Composite T-process regression models. Communications in Mathematics and Statistics, 11(2):307–323, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00249-
                                        4.

Bhushan:2023:NCI

[271] Shashi Bhushan and Abhay Pratap Pandey. New chain imputation methods for estimating population mean in the presence of missing data using two auxiliary variables. Communications in Mathematics and Statistics, 11(2):325–340, June 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-021-00251-w.

Yang:2023:FDE

[272] Jianping Yang, Weiru Chen, and Weiwei Zhuang. Fractional-degree expectation dependence. Communications in Mathematics and Statistics, 11(2):341–368, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00252-
                                        9.

Li:2023:NCI

[273] Junjun Li and Wenquan Cui. A new classifier for imbalanced data based on a generalized density ratio model. Communications in Mathematics and Statistics, 11(2):369–401, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00254-
7.

Liu:2023:SMS

[274] Jingjing Liu, Fang Deng, and Jiansong Deng. Space mapping of spline spaces over hierarchical Tmeshes. Communications in Mathematics and Statistics, 11(2):403–438, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00258-
3.

Qian:2023:TCI

[275] Chao Qian, Zizhou Tang, and Wenjiao Yan. Topology and curvature of isoparametric families in spheres. Communications in Mathematics and Statistics, 11(2):439–475, June 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00259-2.

Bekker:2023:FFM

[276] Andriette Bekker, Farzane Hashemi, and Mohammad Arashi. Flexible factor model for handling missing data in supervised learning. Communications in Mathematics and Statistics, 11(2):477– 501, June 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-021-
00260-9.

Younso:2023:CNN

[277] Ahmad Younso, Ziad Kanaya, and Nour Azhari. Consistency of the k-nearest neighbor classifier for spatially dependent data. Communications in Mathematics and Statistics, 11(3):503–518, September 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s40304-021-
                                        00261-8.

Zhang:2023:NFD

[278] Zebao Zhang. A note on the filtered decomposition theorem. Communications in Mathematics and Statistics, 11(3):519–539, September 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-021-00262-7.

Dai:2023:EGC

[279] Yanmin Dai. Explicit generators of the centre of the quantum group. Communications in Mathematics and Statistics, 11(3):541–562, September 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00263-
                                        6.

Chen:2023:SPA

[280] Zhen-Qing Chen and Yaozhong Hu. Solvability of parabolic Anderson equation with fractional Gaussian noise. Communications in Mathematics and Statistics, 11(3):563–582, September 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00264-
                                        5.

Bao:2023:SSM

[281] Xiaohan Bao, Weihua Tong, and Falai Chen. A spectral segmentation method for large meshes. Communications in Mathematics and Statistics, 11 (3):583–607, September 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00265-4.

Du:2023:RBM

[282] Weitao Du and Elton P. Hsu. Reflecting Brownian motion and the Gauss– Bonnet–Chern theorem. Communications in Mathematics and Statistics, 11(3):609–627, September 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00266-3.

Pu:2023:IEC

[283] Tong Pu, Narayanaswamy Balakrishnan, and Chuancun Yin. An identity for expectations and characteristic function of matrix variate skew-normal distribution with applications to associated stochastic orderings. Communications in Mathematics and Statistics, 11(3):629–647, September 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00267-
2.

Ozen:2023:MSN

[284] Mehmet Özen, Osama A. Naji, Ünsal Tekir, and Suat Koç. On modules satisfying S-Noetherian spectrum condition. Communications in Mathematics and Statistics, 11(3):649–662, September 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00268-
1.

Xiong:2023:NGT

[285] Wenjun Xiong, Juan Ding, Wei Zhang, Aiyi Liu, and Qizhai Li. Nested group testing procedure. Communications in Mathematics and Statistics, 11(4):663–693, December 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-021-00269-0.

Dong:2023:NTR

[286] Liang Dong and Zhishui Hu. The number of triangles in random intersection graphs. Communications in Mathematics and Statistics, 11(4):695–725, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00270-
                                        7.

Ke:2023:APP

[287] Jingyao Ke, Bin Xu, and Zhouwang Yang. Area-preserving parameterization with Tutte regularization. Communications in Mathematics and Statistics, 11(4):727–740, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-021-00271-
                                        6.

Ghoul:2023:PSO

[288] Abdelhak Ghoul, Mokhtar Hafayed, Imad Eddine Lakhdari, and Shahlar Meherrem. Pointwise second–order necessary conditions for stochastic optimal control with jump diffusions. Communications in Mathematics and Statistics, 11(4):741–766, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00272-
5.

Hou:2023:MSN

[289] Songbo Hou. Multiple solutions of a nonlinear biharmonic equation on graphs. Communications in Mathematics and Statistics, 11(4):767–774, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00273-
4.

Li:2023:NAR

[290] Huiqiong Li, Chenchen Ma, Jianguo Sun, and Niansheng Tang. A new approach for regression analysis of multivariate current status data with informative censoring. Communications in Mathematics and Statistics, 11(4):775–794, December 2023. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00274-3.

Yan:2023:SIS

[291] Han Yan and Dehui Wang. Statistical inference for self-exciting threshold INAR processes with missing values. Communications in Mathematics and Statistics, 11(4):795–814, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00275-
2.

Chen:2023:HMW

[292] Qun Chen, Kaipeng Li, and Hongbing Qiu. f-harmonic maps within bounded distance from quasi-isometric maps. Communications in Mathematics and Statistics, 11(4):815–825, December 2023. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00276-
1.

Zhang:2024:WSM

[293] Xicheng Zhang. Weak solutions of McKean–Vlasov SDEs with supercritical drifts. Communications in Mathematics and Statistics, 12(1):1–14, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00277-
0.

Chen:2024:EFD

[294] Chuanjun Chen and Xiaofeng Yang. Efficient fully discrete spectral-Galerkin scheme for the volume-conserved multi– vesicular phase–field model of lipid vesicles with adhesion potential. Communications in Mathematics and Statistics, 12(1):15–43, March 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00278-z.

Shao:2024:SMI

[295] Changguo Shao and Antonio Beltrán. Second maximal invariant subgroups and solubility of finite groups. Communications in Mathematics and Statistics, 12(1):45–54, March 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-021-00279-y.

Ma:2024:TPP

[296] Hong-Yu Ma, Chun-Ming Yuan, and Li-Yong Shen. Tool path planning with confined scallop height error using optimal connected Fermat spirals. Communications in Mathematics and Statistics, 12(1):55–78, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00280-
5.

Huang:2024:LBC

[297] Lihao Huang, Chuanjing Zhang, and Xi Zhang. On the ∂ ¯∂-lemma and Bott– Chern cohomology with local coefficients. Communications in Mathematics and Statistics, 12(1):79–90, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00282-
3.

Sheng:2024:GTT

[298] Mao Sheng and Jinxing Xu. A global Torelli theorem for certain Calabi–Yau threefolds. Communications in Mathematics and Statistics, 12(1):91–112, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-021-00283-
2.

Du:2024:IEG

[299] Jierui Du, Yuan Li, and Xia Cui.

Identification and estimation of generalized additive partial linear models with nonignorable missing response. Communications in Mathematics and Statistics, 12(1):113–156, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00284-
9.

Luo:2024:IRE

[300] Yuxiang Luo, Yang Wei, Zhouping Li, and Bing-Yi Jing. Incorporating relative error criterion to conformal prediction for positive data. Communications in Mathematics and Statistics, 12(1):157–186, March 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-023-00360-
8.

Peng:2024:EBM

[301] Yu Peng, Hao Fu, and Tingsong Du. Estimations of bounds on the multiplicative fractional integral inequalities having exponential kernels. Communications in Mathematics and Statistics, 12(2):187–211, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00285-
8.

Kumar:2024:PGW

[302] Devendra Kumar, Manoj Kumar, and Jagdish Saran. Power generalized Weibull distribution based on record values and associated inferences with bladder cancer data example. Communications in Mathematics and Statistics, 12(2):213–238, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/

article/10.1007/s40304-022-00286-
7.

Duan:2024:PDB

[303] Pingtao Duan, Yuting Liu, and Zhiming Ma. Pricing discrete barrier options under the jump–diffusion model with stochastic volatility and stochastic intensity. Communications in Mathematics and Statistics, 12(2):239–263, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00287-
6.

Wang:2024:RTR

[304] Xiaochen Wang and Xiaomin Zhou. Relative time-restricted sensitivity and entropy. Communications in Mathematics and Statistics, 12(2):265–277, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00289-
4.

Yang:2024:HIE

[305] Fen-Fen Yang. Harnack inequalities for G-SDEs with multiplicative noise. Communications in Mathematics and Statistics, 12(2):279–305, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00290-
x.

Wang:2024:VSD

[306] Haofeng Wang, Xuejun Jiang, Min Zhou, and Jiancheng Jiang. Variable selection for distributed sparse regression under memory constraints. Communications in Mathematics and Statistics, 12(2):307–338, June 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-022-00291-w.

Huang:2024:KEF

[307] Ping Huang, Chenwei Wang, and Ercai Chen. Katok’s entropy formula of unstable metric entropy for partially hyperbolic diffeomorphisms. Communications in Mathematics and Statistics, 12(2):339–355, June 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-022-00293-8.

Zhang:2024:FSP

[308] Li-Xin Zhang. Functional Shige Peng’s central limit theorems for martingale vectors. Communications in Mathematics and Statistics, 12(2):357–383, June 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-022-00294-
                                        7.

Jentzen:2024:SOE

[309] Arnulf Jentzen and Adrian Riekert. Strong overall error analysis for the training of artificial neural networks via random initializations. Communications in Mathematics and Statistics, 12(3):385–434, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-022-00292-
                                        9.

Calcagni:2024:ELL

[310] Antonio Calcagnì. Estimating latent linear correlations from fuzzy frequency tables. Communications in Mathematics and Statistics, 12(3):435–461, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00295-
6.

Lu:2024:SZD

[311] Lu Lu, Lihua Feng, and Weijun Liu.

Signed zero-divisor graphs over commutative rings. Communications in Mathematics and Statistics, 12(3):463–477, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-022-
00297-4.

Cao:2024:FDS

[312] Junying Cao, Jun Zhang, and Xiaofeng Yang. Fully-decoupled and second-order time-accurate scheme for the Cahn–Hilliard Ohta–Kawaski phase-field model of diblock copolymer melt confined in Hele–Shaw cell. Communications in Mathematics and Statistics, 12(3):479–504, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00298-
3.

Yang:2024:DRS

[313] Yi-Jun Yang, Yu-Ming Zhao, Li-Qun Yang, and Wei Zeng. Diffeomorphic registration of 3D surfaces with point and curve landmarks. Communications in Mathematics and Statistics, 12(3):505–522, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00299-
2.

Huang:2024:ACB

[314] Bingru Huang and Falai Chen. An algorithm to compute the μ-bases of rational parametric surfaces with respect to one variable. Communications in Mathematics and Statistics, 12(3):523–541, September 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

//link.springer.com/article/10.
1007/s40304-022-00300-y.

Zhu:2024:DMF

[315] Yujian Zhu and Puying Zhao. Diagnostic measures for functional linear model with nonignorable missing responses. Communications in Mathematics and Statistics, 12(3):543–562, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00301-
x.

Ballester-Bolinches:2024:SCF

[316] A. Ballester-Bolinches, S. F. Kamornikov, V. Pérez-Calabuig, and V. N. Tyutyanov. On σ-solubility criteria for finite groups. Communications in Mathematics and Statistics, 12(3):563–571, September 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00303-

Lin:2024:PEE

[317] Fuming Lin, Yingying Jiang, and Yong Zhou. The k-th power expectile estimation and testing. Communications in Mathematics and Statistics, 12(4):573– 615, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

tronic). URL https://link.springer.
com/article/10.1007/s40304-022-
00302-w.

Ma:2024:FMN

[318] Shilin Ma and Dafeng Zuo. Frobenius manifolds and a new class of extended affine Weyl groups of A-type (II). Communications in Mathematics and Statistics, 12(4):617–632, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00305-
7.

Deng:2024:IAI

[319] Fang Deng, Tianhui Yang, Jingjing Liu, and Jiansong Deng. Isogeometric analysis on implicit domains: Approximation, stability and error estimates. Communications in Mathematics and Statistics, 12(4):633–658, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00307-
5.

Wei:2024:ADR

[320] Shaojie Wei, Gaorong Li, and Zhongzhan Zhang. An alternative doubly robust estimation in causal inference model. Communications in Mathematics and Statistics, 12(4):659–678, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-022-00308-
                                        4.

Fakhfakh:2024:CSK

[321] Raouf Fakhfakh and Abdelhamid Hassairi. Cauchy–Stieltjes kernel families and free multiplicative convolution. Communications in Mathematics and Statistics, 12(4):679–694, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-022-00311-
                                        9.

Eidi:2024:FHG

[322] Marzieh Eidi and Jürgen Jost. Floer homology: From generalized Morse– Smale dynamical systems to Forman’s combinatorial vector fields. Communications in Mathematics and Statistics, 12(4):695–720, ???? 2024. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s40304-022-00314-6.

Yang:2024:DPF

[323] Quan-Hui Yang and Lilu Zhao. On a discriminator for the polynomial f(x) = x^{3} + x. Communications in Mathematics and Statistics, 12(4):721–734, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s40304-022-00315-
                                        5.

Xu:2024:ECS

[324] Jinyan Xu and Liang Zhao. Existence and convergence of solutions for nonlinear elliptic systems on graphs. Communications in Mathematics and Statistics, 12(4):735–754, ???? 2024. CODEN ???? ISSN 2194-6701 (print), 2194-671X (electronic).

URL https://link.springer.com/
article/10.1007/s40304-022-00318-